### v1

# Active Search for Low-altitude UAV Sensing and Communication for Users at Unknown Locations

Abstract
This paper studies optimal unmanned aerial vehicle (UAV) placement to ensure line-of-sight (LOS) communication and sensing for a cluster of ground users possibly in deep shadow, while the UAV maintains backhaul connectivity with a base station (BS). The key challenges include unknown user locations, uncertain channel model parameters, and unavailable urban structure. Addressing these challenges, this paper focuses on developing an efficient online search strategy which jointly estimates channels, guides UAV positioning, and optimizes resource allocation. Analytically exploiting the geometric properties of the equipotential surface, this paper develops an LOS discovery trajectory on the equipotential surface while the closed-form search directions are determined using perturbation theory. Since the explicit expression of the equipotential surface is not available, this paper proposes to locally construct a channel model for each user in the LOS regime utilizing polynomial regression without depending on user locations or propagation distance. A class of spiral trajectories to simultaneously construct the LOS channels and search on the equipotential surface is developed. An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel. Numerical results on real 3D city maps demonstrate that the proposed scheme achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search.
Index Terms:
UAV trajectory design, unknown user locations, unknown channels, unknown environment, perturbation theory
I Introduction
In dense urban environments, short-wavelength signals such as millimeter wave (mmW), Terahertz (THz), and optical waves experience significant attenuation due to obstructions like buildings and other structures [1]. This attenuation is particularly problematic for sensing performance, which depends heavily on maintaining line-of-sight (LOS) links between targets and transceivers [2, 3].
Deploying a low-altitude unmanned aerial vehicle (UAV) presents a promising solution for establishing LOS links for users in deep shadowed areas, as UAV s operate at relatively high altitudes with the flexibility of 3D mobility. However, optimizing UAV deployment remains challenging due to the arbitrarily complex terrain topology, which complicates modeling and predicting LOS conditions. Previous works have circumvented this challenge by assuming distance-dependent pure LOS channel models [4, 5], or probabilistic channel models [6, 7, 8, 9]. While these simplified models allow for coarse analysis, they often compromise performance by neglecting the actual LOS conditions. Some state-of-the-art works have developed blockage-aware channel models [10, 11] to enhance UAV placement strategies. For instance, the authors in [10] approximated buildings as polyhedrons and derived explicit expressions to determine the LOS status of a 3D position using an offline city map. However, constructing geometric maps or radio maps incurs substantial costs, and polyhedral approximation, while useful, remains a coarse and computationally intensive method, making it challenging to scale to larger areas.
In addition, most existing studies [4, 5, 6, 7, 8, 9, 10, 11, 12] assume perfect knowledge of user locations and channel models. However, in practical scenarios, accurate user locations or precise channel models are typically unavailable due to the location privacy of the users and the uncertain non-line-of-sight (NLOS) condition. Some recent research efforts have considered first estimating user locations and channel parameters before optimizing UAV service positions [13, 14, 15]. Nevertheless, localization under potential NLOS conditions is already a challenging problem, and the performance of UAV placement is highly dependent on the accuracy of user localization. Assuming the availability of 3D maps, the work in [15] employed the Fisher information metric to design an online trajectory for joint path-loss parameter estimation and user localization, although the global optimality and complexity of the approach remain unclear. With known user locations and maps, the authors in [13] utilized dynamic programming to optimize UAV trajectory for channel parameter learning, employing map compression to estimate LOS probability and assist in UAV placement. However, the effectiveness of the placement heavily depends on channel learning and LOS probability estimation. In the absence of prior knowledge of channels and user locations, a model-free approach presented in [14] relied on signal-to-interference-and-noise ratio (SINR) measurements and a 3D map to account for blockages and scatterings, applying Q-learning to optimize UAV placement. Despite this, the action space was limited to four orthogonal horizontal directions, significantly reducing the likelihood of identifying the optimal UAV placement in a 3D environment.
This paper investigates the UAV trajectory design to provide communication and sensing services for a cluster of ground users at unknown locations, while also maintaining backhaul connectivity with a remote base station (BS). The goal is to develop an optimization framework for an efficient online search of the best UAV position for a broad class of communication problems and sensing problems. The challenges arise from the need to operate without precise information: user locations are unknown due to privacy concerns and potential NLOS conditions, the channel model parameters are difficult to determine due to environmental sensitivity, and the lack of accurate 3D maps further complicates the UAV placement in dense urban environments.
In this work, we develop three key techniques to address the aforementioned challenges. First, we introduce a search method on an equipotential surface using trajectories derived from perturbation theory. The equipotential surface represents a region where the sensing and communication performance for users is balanced with the capacity of the UAV-BS backhaul link. The UAV’s proximity to the equipotential surface can be assessed by measuring the signal-to-noise ratio (SNR), without requiring knowledge of user locations or channel parameters. Although the exact expression of the equipotential surface remains elusive, perturbation theory guides the UAV search direction while maintaining its position on this surface. Second, we propose a local construction of channel models for each user within the LOS regime using local polynomial regression. This approach allows for a relatively simple construction of local LOS models without requiring user location or propagation distance. Third, we develop a search strategy on the equipotential surface that leverages two universal properties: upward invariance and colinear invariance of LOS regions over almost any terrain structure. Prior studies have shown that, in certain cases, this strategy can achieve an $\epsilon$-optimal solution globally in 3D space with a trajectory length proportional to the search radius [16, 17]. However, previous work [17] assumed knowledge of user locations and channel model parameters.
Contributions of this paper are as follows:
  • We study UAV trajectory design for serving a cluster of users at unknown locations with unknown channels. Specifically, we develop our search strategy based on perturbation theory and two universal properties, i.e., upward invariance and colinear invariance, of the LOS regions over almost all terrain structure.
  • We provide an optimal design for a class of UAV search trajectories to balance the estimation for user channels and the exploration for better channel quality. This also includes deriving the upper bound of the mean squared error (MSE) for the locally constructed channel gain based on an optimal measurement pattern and determining the optimal radius of the spiral trajectory for minimizing the MSE of channel gain estimation.
  • We demonstrate through numerical results that the normalized channel gain construction error is on the order of $10^{-2}$ without knowing user locations or propagation distances. Furthermore, with a $3$ kilometer search, the proposed scheme can achieve over $95\%$ performance of that from a 3D exhaustive search over a dense urban area.
The remaining part of the paper is organized as follows: Section II introduces the system model and formulates a UAV-assisted sensing and communication problem. Section III discusses the geometric properties of the equipotential surface and local channel model construction. Section IV outlines the LOS discovery trajectory and proposes a superposed trajectory for optimal LOS position search. Section V presents numerical results and comparisons, while Section VI concludes the paper.
Notation: Vectors and matrices are denoted by bold $\mathbf{x}$ and bold capital $\mathbf{X}$, respectively. $\mathcal{M}_{m,n}$ denotes all $m$-by-$n$ matrices with $\mathcal{M}_{n}$ for square matrices. Matrix entry, column vector, and row vector are represented as $[\mathbf{X}]_{(i,j)}$, $[\mathbf{X}]_{(:,j)}$, and $[\mathbf{X}]_{(i,:)}$, respectively. Matrix trace and diagonal are $\text{tr}\{\mathbf{X}\}$ and $\text{diag}\{\mathbf{X}\}$. Expectation and variance are $\mathbb{E}\{\cdot\}$ and $\mathbb{V}\{\cdot\}$. The gradient of $f(\mathbf{x})$ is $\nabla f(\mathbf{x})$. The cross product is indicated by $\times$. The time derivative of $x(t)$ is defined as $\dot{x}=\text{d}x(t)/\text{d}t$. $C$ represents a constant. The inequality $\mathbf{p}\geq 0$ indicates that all the entries in $\mathbf{p}$ are no less than 00.
II System Model
In this section, we first establish an environment model with certain LOS properties. Then, a channel model for LOS propagation condition and NLOS propagation condition is defined. Finally, a general LOS-guaranteed UAV-assisted sensing and communication problem is formulated with two specific problems followed.
II-A Environment Model
Consider that a UAV serves one BS located at $\mathbf{u}_{0}$ and a cluster of users without knowing their locations as shown in Fig. 1. We focus on the case where the users are clustered in an unknown neighborhood in a dense urban area. The sets of users are denoted as $\mathcal{K}\triangleq\{1,2,\dots,K\}$.
Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{3}:x_{3}\geq H_{\text{min}}\}$ be the feasible region of the UAV positions. While signals of ground users are likely blocked by buildings, we denote $\mathcal{D}_{k}\text{$\subseteq$}\mathcal{X}$ as the region of UAV positions such that there is an LOS link between the UAV at position $\mathbf{x}\in\mathcal{D}_{k}$ and user $k$ at an unknown position $\mathbf{u}_{k}$.
The LOS region $\mathcal{D}_{k}$ can be arbitrary except that $\mathcal{D}_{k}$ is assumed to have the following properties: For any $\mathbf{x}\in\mathcal{D}_{k}$,
  1. 1.
    Upward invariant: any UAV position $\mathbf{x}^{\prime}$ perpendicularly above $\mathbf{x}$ also belongs to $\mathcal{D}_{k}$, i.e., $\mathbf{x}^{\prime}\in\mathcal{D}_{k}$;
  2. 2.
    Colinear invariant: any UAV position $\mathbf{x}^{\prime}$ that satisfies $\mathbf{x}^{\prime}-\mathbf{u}_{k}=\rho(\mathbf{x}-\mathbf{u}_{k})$ for some $\rho>1$ also belongs to $\mathcal{D}_{k}$.
Similarly, one can define $\mathcal{D}_{0}\subseteq\mathcal{X}$ as the LOS region of UAV positions to the BS. To summarize, the upward invariant and colinear invariant properties imply that if there is an LOS link between the UAV and a user, such an LOS condition will remain if the UAV increases its altitude or moves away from the user without changing the elevation and azimuth angles. The widely adopted probabilistic LOS model in the UAV literature [18, 19, 20, 21] is a special case that satisfies these properties in a statistical sense.
Define the full-LOS region $\tilde{\mathcal{D}}=\bigcap\mathcal{D}_{k}$ as the set of UAV positions where there are LOS links to the BS and all the users. Since the full-LOS region is an intersection of $\mathcal{D}_{k}$, the upward invariant property automatically holds, i.e., for any full-LOS position $\mathbf{x}\in\tilde{\mathcal{D}}$, any position $\mathbf{x}^{\prime}$ perpendicularly above $\mathbf{x}$ is also a full-LOS position which satisfies $\mathbf{x}^{\prime}\in\tilde{\mathcal{D}}$. Note that the colinear invariant property does not hold for $\tilde{\mathcal{D}}$.
II-B Channel Model
Based on the LOS region $\mathcal{D}_{k}$, the channel gain from the UAV at position $\mathbf{x}$ to node $k$ at the unknown position $\mathbf{u}_{k}$ is modeled as
$\bar{g}_{k}(\mathbf{x})=\begin{cases}\begin{aligned} &g_{k}(\mathbf{x})&&\text{if }\mathbf{x}\in\mathcal{D}_{k}\quad\text{(1a)}\\ &g_{k}(\mathbf{x})+\phi(\mathbf{x})&&\text{otherwise}\quad\,\text{(1b)}\end{aligned}\end{cases}$ (1)
where $g_{k}(\mathbf{x})$ is the deterministic channel gain under LOS, and $\phi(\mathbf{x})$ is a random variable to capture the power penalty due to the shadowing in NLOS condition [22].
The global model for the deterministic channel gain $g_{k}(\mathbf{x})$ is unknown to the system, except that $g_{k}(\mathbf{x})$ is assumed to be Lipschitz continuous satisfying,
$g_{k}(\mathbf{x})$ $\leq g_{k}(\mathbf{x}_{0})+\nabla g_{k}(\mathbf{x}_{0})^{\text{T}}(\mathbf{x}-\mathbf{x}_{0})+\frac{L_{g}}{2}d^{2}(\mathbf{x},\mathbf{x}_{0})$ (2)
and
$g_{k}(\mathbf{x})$ $\geq g_{k}(\mathbf{x}_{0})+\nabla g_{k}(\mathbf{x}_{0})^{\text{T}}(\mathbf{x}-\mathbf{x}_{0})-\frac{L_{g}}{2}d^{2}(\mathbf{x},\mathbf{x}_{0})$ (3)
for all $\mathbf{x}\in\mathcal{X}$ where $L_{g}$ is a finite constant, [^1] The free-space propagation model, i.e., $g_{k}(\mathbf{x})=b_{0}+10a_{0}\log_{10}(d(\mathbf{x},\mathbf{u}_{k}))$ where $b_{0}$ and $a_{0}$ are channel parameters, satisfies this condition with $L_{g}=8.7\times\text{10}^{-4}$ under $a_{0}=-2$ and $H_{\text{min}}=100$ meters. and $d(\mathbf{x},\mathbf{x}_{0})\triangleq\|\mathbf{x}-\mathbf{x}_{0}\|_{2}$.
While the global model $g_{k}({\bf x})$ or $\bar{g}_{k}({\bf x})$ is not available, the UAV may measure the channel gain $\bar{g}_{k}(\mathbf{x})$ when it explores location $\mathbf{x}$. The measurement model is given by
$y=\bar{g}_{k}(\mathbf{x})+\xi$ (4)
where $\xi\sim\text{$\mathcal{N}$}(0,\sigma^{2})$ models the small-scale fading and the uncertainty due to the antenna gain which may not be omidirectional.
In addition, the complete geometry of $\mathcal{D}_{k}$ is also unknown, except for a local area that has been explored by the UAV along its trajectory up to time $t$. Specifically, denote $\mathbf{x}(t)$ as the UAV position at time $t$. We assume that the value of LOS indicator function $\mathbb{I}\{\mathbf{x}(t)\in\mathcal{D}_{k}\}$ can be perfectly determined based on the measurements $\bar{g}_{k}(\mathbf{x}(\tau))$ for $0\leq\tau\leq t$. For a practical implementation, $\mathbb{I}\{\mathbf{x}(t)\in\mathcal{D}_{k}\}$ can be computed by statistical learning and hypothesis testing [22, 12].
Figure 1: Illustration of a system where a UAV provides sensing and/or communication services for a cluster of sensing targets and/or communication users without knowing their locations, channel models, and the city topology while establishing an LOS backhaul link to a BS.
II-C UAV-assisted Sensing and Communication with Backhaul
A common joint UAV position optimization and resource allocation problem needs to balance the performance of serving the users and the capacity of the backhaul link that the UAV connects to the BS. In addition, most sensing tasks require an LOS condition. Likewise, for a communication task, the LOS condition can substantially enhance the communication performance. These requirements lead to a general max-min problem for the LOS-guaranteed UAV-assisted sensing and communication as follows.
Denote $\bm{g}_{\text{u}}(\mathbf{x})$ as a vector that collects the LOS channel gains for all the users, i.e., $\bm{g}_{\text{u}}(\mathbf{x})=[g_{1}(\mathbf{x}),\dots,g_{K}(\mathbf{x})]^{\text{T}}$. Denote $\mathbf{p}$ as the corresponding resource allocation given $\bm{g}_{\text{u}}(\mathbf{x})$. We have
$\mathscr{P}:\quad\mathop{\mbox{maximize}}\limits_{\mathbf{x},\mathbf{p}}$ $\quad\min\{f_{0}(g_{0}(\mathbf{x})),F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\}$ (5)
[…]
$\quad H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0,n=1,2,\dots,N$
where $f_{0}(g_{0}(\mathbf{x}))$ represents the objective of the BS-UAV link under the LOS condition, $F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ represents the objective of the UAV-user links under LOS, and $H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0$ are the corresponding constraints for the resource allocation. Problem 5 is non-convex due to the LOS constraint on UAV positions and the fact that the blockage can have an arbitrary shape.
Such a general formulation captures many typical applications for communication and sensing, with two examples illustrated as follows.
II-C1 Balancing problem
Consider to deploy a UAV to offer sensing and communication services for a cluster of sensing targets and communication users while establishing a LOS relay link with a remote BS. Specify $p_{k}$ as the power allocation from the UAV to user $k$. Denote $f_{k}(g_{k}(\mathbf{x}),p_{k})$ as the objective function for user $k$, and $N_{0}$ as the noise power of the propagation channel. For sake of brevity, let $N_{0}=1$ in this paper. For a sensing task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as weighted SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{s}}p_{k}g_{k}(\mathbf{x})$ where $\mu_{\text{s}}$ is a weight of the sensing task [23, 24]. For a communication task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as channel capacity, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{c}}\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$ where $\mu_{\text{c}}$ is a weight of the communication task. In addition, the objective function of the BS-UAV link is given by the capacity function, i.e., $f_{0}(g_{0}(\mathbf{x}))=\log_{2}(1+P_{0}g_{0}(\mathbf{x}))$ where $P_{0}$ is the transmit power of the BS. The balancing problem aims at maximizing the worst link performance of the BS-UAV link and the UAV-user links. Hence, the overall objective function is specified as $\min\{f_{0}(g_{0}(\mathbf{x})),\min_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k})\}\}$. The balancing problem jointly optimizes the UAV position $\mathbf{x}$ and the power allocation $\mathbf{p}$ as follows
$\mathop{\mbox{maximize}}\limits_{\mathbf{x},\mathbf{p}\geq 0}$ $\quad\min\{f_{0}(g_{0}(\mathbf{x})),\min_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k}\text{)}\}\}$ (6)
[…]
where $P_{\text{T}}$ is the total transmit power of the UAV.
II-C2 Sum-rate problem
Consider that a UAV relays signal from a BS to $K$ ground users with other assumptions the same as that in the balancing problem. The sum-rate problem aims at maximizing the sum capacity of the relay channels. Thus, the objective function is $\min\{f_{0}(g_{0}(\mathbf{x})),\sum_{k\in\mathcal{K}}f_{k}(g_{k}(\mathbf{x}),p_{k})\}$ where $f_{0}(g_{0}(\mathbf{x}))$ and $f_{k}(g_{k}(\mathbf{x}),p_{k})$ are defined in Section II-C1, and the problem is formulated as
$\mathop{\mbox{maximize}}\limits_{\mathbf{x},\mathbf{p}\text{$\geq$}0}$ $\quad\min\{f_{0}(g_{0}(\mathbf{x})),{\textstyle\sum_{k\in\mathcal{K}}}f_{k}(g_{k}(\mathbf{x}),p_{k})\}$ (7)
[…]
$\quad\sum_{k\in\mathcal{K}}p_{k}\leq P_{\text{T}}.$
While the balancing problem is non-convex and the sum-rate problem is convex in the resource allocation variable $\mathbf{p}$, the whole problem $\mathscr{P}$ that jointly optimizes the UAV location $\mathbf{x}$ is non-convex. In addition, as the channels and the LOS regions $\mathcal{D}_{k}$ are not available before exploring near location $\mathbf{x}$, the UAV needs to design an online trajectory to explore the LOS opportunity, measure the channel quality, and optimize for the system performance.
II-D Suboptimal Solution on the Equipotential Surface
As the full LOS region $\tilde{\mathcal{D}}$ can have an arbitrary shape and is initially unknown before the exploration, finding the globally optimal solution to $\mathscr{P}$ generally requires an online exhaustive search in 3D, which is prohibitive due to the limited flight time of UAV. Thus, we compromise for a suboptimal solution on the equipotential surface.
The equipotential surface $\mathcal{S}$ is defined as a region where the objective of the BS-UAV link and the objective of the UAV-user links under the optimized resource allocation are equal, assuming all the links were in LOS. Specifically, define $\mathbf{p}^{*}(\mathbf{x})$ as the optimal solution to $\mathscr{P}$ given a fixed location $\mathbf{x}$ by ignoring the LOS constraint $\mathbf{x}\in\text{$\tilde{\mathcal{D}}$}$. Denote $\bm{g}(\mathbf{x})=[g_{0}(\mathbf{x}),\bm{g}_{\text{u}}(\mathbf{x})^{\text{T}}]^{\text{T}}$ as a vector that collects the LOS channel gains from the BS and the users. Defining a balance function
$F(\bm{g}(\mathbf{x}))\text{$\triangleq$}f_{0}(g_{0}(\mathbf{x}))-F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p}^{*}(\mathbf{x}))$ (8)
the equipotential surface is defined as
$\mathcal{S}=\left\{\mathbf{x}\in\mathcal{X}:F(\bm{g}(\mathbf{x}))=0\right\}.$ (9)
Recent studies [17, 12, 16] discover that searching on the equipotential surface $\mathcal{S}$ has significant promise in identifying the globally optimal UAV position in 3D space. It has been shown that for the case with a single user and a BS, a solution can be found with only an $O(\epsilon)$ performance gap to the globally optimal solution by searching only on the equipotential surface, for a search distance of $O(1/\epsilon)$, where the equipotential surface degenerates to a middle perpendicular plane under the condition that $P_{0}=P_{\text{T}}$ [12, 16]. It was also numerically demonstrated in [17], where the user locations are known, that searching on the equipotential surface in a multi-user case for a sum-rate maximization objective attains over $96\%$ of the performance of that from an exhaustive search in the entire 3D space.
II-E Superposed Trajectory for Unknown User Locations
Since the user locations $\mathbf{u}_{k}$ are unknown, the analytical form of the channels $\bm{g}(\mathbf{x})$ is not available, and hence, no analytical form of the equipotential surface $\mathcal{S}$ is available to the system. As a result, while the UAV aims at searching on $\mathcal{S}$, it also needs to simultaneously estimate $\bm{g}(\mathbf{x})$ and construct $\mathcal{S}$. A classical approach may first estimate a parametric form of $\bm{g}(\mathbf{x})$ and then construct a global analytical model for $\mathcal{S}$. However, it is known that a joint user localization and propagation parameter estimation for $\bm{g}(\mathbf{x})$ require measurements across a large area, and such global construction is very challenging and inaccurate.
To tackle this challenge, we develop a superposed trajectory for the UAV exploration as $\mathbf{x}(t)=\text{{x}}_{\text{s}}(t)+\text{{x}}_{\text{r}}(t)$ where $\text{{x}}_{\text{s}}(t)$ is the online search trajectory to find a suboptimal solution to $\mathscr{P}$ on $\mathcal{S}$ based on the local information of $\bm{g}(\mathbf{x})$ in the neighborhood along the search, and $\text{{x}}_{\text{r}}(t)$ provides small deviation from $\text{{x}}_{\text{s}}(t)$ to simultaneously collect measurements for the local reconstruction of $\bm{g}(\mathbf{x})$ and $\mathcal{S}$.
III Trajectory Design for Tracking on the Equipotential Surface
In this section, we first explore the geographic properties of $\mathcal{S}$, and subsequently, demonstrate the feasibility of approximately constructing the equipotential surface without knowing the channel models and the user locations. Then, the local construction of the propagation model is studied with theoretical results on the optimal measurement pattern. A class of spiral trajectories for locally constructing the equipotential surface is proposed at the end of this section.
III-A Property of the Equipotential Surface
The equipotential surface does not exist when $\mathcal{S}=\text{$\varnothing$}$. This corresponds to a superior channel for the BS or for the users, resulting in a trivial solution where the UAV should hover above either the BS or above the cluster of the users. We focus on the scenario where $\mathcal{S}$ does exist.
III-A1 Existence Condition
The existence of the equipotential surface can be easily checked by evaluating the objective values at two special locations $\mathbf{x}_{0}^{\text{m}}=\mathbf{u}_{0}+[0,0,H_{\text{min}}]^{\text{T}}$ and $\mathbf{x}_{\text{u}}^{\text{m}}=\sum_{k\in\mathcal{K}}\mathbf{u}_{k}/K+[0,0,H_{\text{min}}]^{\text{T}}$. A general existence condition yields $F(\bm{g}(\mathbf{x}_{0}^{\text{m}}))F(\bm{g}(\mathbf{x}_{\text{\text{u}}}^{\text{m}}))\leq 0$. This is because both the channel $\bm{g}(\mathbf{x})$ and $F(\bm{g}(\mathbf{x}))$ are continuous, and hence, there exists a path from $\mathbf{x}_{\text{u}}^{\text{m}}$ to $\mathbf{x}_{0}^{\text{m}}$ that reaches $F(\bm{g}(\mathbf{x}))=0$.
For a specific problem, such as the balancing problem in (6), the existence condition depends on the problem parameters, such as the power budget $P_{\text{T}}$.
Proposition 1 (Existence condition in a specified balancing problem).
For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, a sufficient condition to the existence of the equipotential surface is given by
$\left(\frac{P_{0}}{P_{{\rm T}}}-\frac{1/g_{0}(\mathbf{x}_{0}^{\text{{\rm m}}})}{\sum_{k\in\mathcal{K}}(1/g_{k}(\mathbf{x}_{0}^{{\rm m}}))}\right)\left(\frac{P_{0}}{P_{{\rm T}}}-\frac{1/g_{0}({\rm\mathbf{x}_{\text{{\rm u}}}^{{\rm m}}})}{\sum_{k\in\mathcal{K}}(1/g_{k}({\rm\mathbf{x}_{{\rm u}}^{{\rm m}}}))}\right)\leq 0.$ (10)
Proof.
See Appendix A. ∎
Proposition 1 provides an explicit condition to the existence of the equipotential surface in a typical balancing problem for maximizing the worst relay channel capacity. In particular, when $\mathcal{K}=\{1\}$, $1/(g_{0}(\mathbf{x})\sum_{k\in\mathcal{K}}(1/g_{k}(\mathbf{x})))$ is simplified as $g_{1}(\mathbf{x})/g_{0}(\mathbf{x})$, and thus the sufficient condition in (10) becomes $(P_{0}/P_{\text{T}}-g_{1}(\mathbf{x}_{0}^{\text{m}})/g_{0}(\mathbf{x}_{0}^{\text{m}}))(P_{0}/P_{\text{T}}-g_{1}(\mathbf{x}_{\text{u}}^{\text{m}})/g_{0}(\mathbf{x}_{\text{u}}^{\text{m}}))\leq 0$. Consequently, there exists a point $\mathbf{x}$ satisfying $P_{0}/P_{\text{T}}=g_{1}(\mathbf{x})/g_{0}(\mathbf{x})$ implying that a smaller channel gain necessitates a corresponding increment in the allocated power budget.
III-A2 Geometric Shape
While the geometric characteristics of the equipotential surface $\mathcal{S}$ are highly related to the specific applications, the exploitation of geometric properties of $\mathcal{S}$ in a typical balancing problem provides some insights for UAV trajectory design on $\mathcal{S}$.
Proposition 2 (A spherical equipotential surface).
Suppose that the channel gain satisfies $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$. For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the equipotential surface is a sphere centered at $\mathbf{o}=(P_{0}\sum_{k\in\mathcal{K}}\mathbf{u}_{k}-P_{\text{T}}\mathbf{u}_{0})/(KP_{0}-P_{\text{T}})$ with radius $R$ satisfying
$R^{2}=\frac{P_{\text{T}}\|\mathbf{u}_{0}\|_{2}^{2}-P_{0}\sum_{k\in\mathcal{K}}\|\mathbf{u}_{k}\|_{2}^{2}}{KP_{0}-P_{\text{T}}}+\|\mathbf{o}\|_{2}^{2}$ (11)
Proof.
See Appendix B. ∎
Knowing that the equipotential surface $\mathcal{S}$ is a sphere, one may estimate the center and radius of the sphere, and hence, one can easily design a trajectory exploring LOS opportunity on $\mathcal{S}$. In general, when $\mathcal{S}$ is not guaranteed to be a sphere, it is also inspired from Proposition 2 that $\mathcal{S}$ may be locally approximated by a patch from a sphere, leading to some simplified design of local trajectories. In particular, if $K=1$ and $P_{\text{T}}=P_{0}$, the equipotential surface becomes a middle-perpendicular plane between the BS and the user satisfying $d(\mathbf{x},\mathbf{u}_{0})=d(\mathbf{x},\mathbf{u}_{1})$ for $\forall\mathbf{x}\in\mathcal{S}$ [16].
III-B Trajectory towards the Equipotential Surface
Recall that the channel gain $\bm{g}(\mathbf{x})$ is not available before the UAV explores location $\mathbf{x}$, and moreover, the analytical form of $F(\bm{g}(\mathbf{x}))$ in (8) is not available. Here, we develop an iterative exploration strategy to move towards the equipotential surface $\mathcal{S}$ that satisfies $F(\bm{g}(\mathbf{x}))=0$.
We adopt a gradient-type search. Whenever the UAV locates off the equipotential surface $F(\bm{g}(\mathbf{x}))=\delta\neq 0$ at $\mathbf{x}=\mathbf{c}_{0}$, it moves to the direction that steepest decreases (or increases) $F(\bm{g}(\mathbf{x}))$. A linear approximation of $F(\bm{g}(\mathbf{x}))$ at $\mathbf{x}=\mathbf{c}_{0}$ yields
$F(\bm{g}(\mathbf{x}))\approx F(\bm{g}(\mathbf{c}_{0}))+\nabla F(\bm{g}(\mathbf{c}_{0}))^{\text{T}}\mathbf{G}(\mathbf{c}_{0})(\mathbf{x}-\mathbf{c}_{0})$ (12)
where $\nabla F(\bm{g}(\mathbf{c}_{0}))^{\text{T}}\mathbf{G}(\mathbf{c}_{0})$ represents the gradient of $F(\bm{g}(\mathbf{x}))$ at $\mathbf{x}=\mathbf{c}_{0}$ and $\mathbf{G}(\mathbf{x})=[\nabla g_{0}(\mathbf{x}),\nabla g_{1}(\mathbf{x}),\dots,\nabla g_{K}(\mathbf{x})]^{\text{T}}$ is the matrix that collects the gradients of the channel gains $\bm{g}(\mathbf{x})$.
Setting the above linear approximation (12) to 00 and noticing that $F(\bm{g}(\mathbf{c}_{0}))=\delta$, a nearest solution $\mathbf{x}$ from $\mathbf{c}_{0}$ can be found by
$\mathop{\mbox{minimize}}\limits_{\mathbf{x}}$ $\quad\|\mathbf{x}-\mathbf{c}_{0}\|_{2}^{2}$ (13)
[…]
where the closed-form solution is obtained using the Lagrangian multiplier method as $\hat{\mathbf{x}}=\mathbf{c}_{0}+\mathcal{V}(\mathbf{c}_{0})$, where
$\mathcal{V}(\mathbf{c}_{0})\triangleq-\frac{\mathbf{G}(\mathbf{c}_{0})^{\text{T}}\nabla F(\bm{g}(\mathbf{c}_{0}))F(\bm{g}(\mathbf{c}_{0}))}{\|\mathbf{G}(\mathbf{c}_{0})^{\text{T}}\nabla F(\bm{g}(\mathbf{c}_{0}))\|_{2}^{2}}.$ (14)
The solution $\hat{\mathbf{x}}$ provides an estimated location on $\mathcal{S}$. Consequently, the one-step exploration range for the UAV to approximately reach $\mathcal{S}$ is given by $r_{0}\triangleq\|\hat{\mathbf{x}}-\mathbf{c}_{0}\|_{2}$.
As a result, a search trajectory for tracking $\mathcal{S}$ can be designed as $\mathbf{x}_{\text{s}}(t_{n+1})=\mathbf{x}_{\text{s}}(t_{n})+\mathcal{V}(\mathbf{x}_{\text{s}}(t_{n}))$. For mathematical convenience, the search trajectory $\mathbf{x}_{\text{s}}(t)$ that is described by a continuous-time ordinary differential equation (ODE) is given by
$\dot{\mathbf{x}}_{\text{s}}=\text{d}\mathbf{x}_{\text{s}}(t)/\text{d}t=\mu_{\text{v}}\mathcal{V}(\mathbf{x}_{\text{s}}(t))$ (15)
where $\mu_{\text{v}}>0$ is a step size to control the speed of the trajectory.
It is observed that computing the search direction $\mathcal{V}(\mathbf{x})$ in (14) not only requires the channel gains $\bm{g}(\mathbf{x})$, but also its gradient $\mathbf{G}(\mathbf{x})$. In the rest of this section, we develop methods and trajectory $\mathbf{x}_{\text{r}}(t)$ to locally construct $\bm{g}(\mathbf{x})$ and its gradient $\mathbf{G}(\mathbf{x})$ at the neighborhood of $\mathbf{x}_{\text{s}}(t)$.
III-C Construction of a Local Channel Model
It is well-known that directly estimating a global propagation model $g_{k}(\mathbf{x})$ is difficult, as it leads to nonlinear regression. Instead, for each node $k$, we adopt a linear model $\hat{g}(\mathbf{x})$ to locally approximate the LOS channel gain $g_{k}(\mathbf{x})$ for the LOS region $\mathbf{x}\in\mathcal{D}_{k}$ in the neighborhood of $\mathbf{x}=\mathbf{c}_{0}$:
$\hat{g}(\mathbf{x})=\alpha+\boldsymbol{\beta}^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})$ (16)
where $\mathbf{\boldsymbol{\theta}}\triangleq[\alpha,\boldsymbol{\beta}^{\text{T}}]^{\text{T}}$, in which, $\alpha\in\mathbb{R}$ and $\boldsymbol{\beta}=[\beta_{1},\beta_{2},\beta_{3}]^{\text{T}}\in\mathbb{R}^{3}$ are channel parameters to be estimated based on the LOS measurements modeled in (4).
For each node $k$, let $\{(\mathbf{x}_{m},y_{m}),m=1,2,\dots,M\}$ be the set of measurements where all the measurements are assumed taken in the LOS case and $y_{m}$ is the noisy measurement (4) with respect to (w.r.t.) to user $k$ at $\mathbf{x}_{m}=\mathbf{x}(t_{m})$, where the measurement noise $\xi_{m}$ is assumed to be independent. Note that one can easily differentiate LOS from NLOS measurements by simply tracking the received signal strength. A least-squares solution of $\mathbf{\boldsymbol{\theta}}$ can be derived as [25, 26]
$\hat{\boldsymbol{\theta}}\triangleq\left[\begin{array}[]{c}\hat{\alpha}\\ \hat{\boldsymbol{\beta}}\end{array}\right]=\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\mathbf{y}$ (17)
where $\mathbf{y}=[y_{1},y_{2},\dots,y_{M}]^{\text{T}}$, and $\tilde{\mathbf{X}}=[\boldsymbol{1},\mathbf{X}]$ in which, $\boldsymbol{1}$ is a column vector of all $1$s, and $\mathbf{\mathbf{X}}$ is an $M\times 3$ matrix with the $m$th row given by $(\mathbf{x}_{m}-\mathbf{c}_{0})^{\text{T}}$.
As a result, for each user $k$, the channel gain and gradient at $\mathbf{x}=\mathbf{c}_{0}$ for computing $\mathcal{V}(\mathbf{x})$ in (14) are estimated as $g_{k}(\mathbf{c}_{0})=\hat{\alpha}$ and $\nabla g_{k}(\mathbf{c}_{0})=\hat{\bm{\beta}}$.
III-D Optimal Measurement Pattern
From the least-squares solution $\hat{\boldsymbol{\theta}}$ in (17), the measurement pattern $\tilde{\boldsymbol{\mathbf{X}}}$ affects the construction performance. We optimize the measurement pattern $\tilde{\boldsymbol{\mathbf{X}}}$ via analyzing the estimation error $\boldsymbol{\theta}^{\text{(e)}}=\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$.
Theorem 1 (Minimum variance).
Given $d(\mathbf{x}_{m},\mathbf{c}_{0})\leq r_{1}$ for all $m=1,2,\dots,M$, the lower bound of the variance of the estimation error $\boldsymbol{\theta}^{\text{{\rm(e)}}}$ is given by
$\text{{\rm tr}}\left\{\mathbb{V}\{\boldsymbol{\theta}^{\text{{\rm(e)}}}\}\right\}\geq\frac{\sigma^{2}}{M}+\frac{9\sigma^{2}}{Mr_{1}^{2}}$ (18)
with equality achieved when the following conditions are satisfied for all coordinates $j,j^{\prime}\in\{1,2,3\}$: (i)$\sum_{m=1}^{M}(x_{mj}-c_{0j})=0$; (ii) $\sum_{m=1}^{M}(x_{mj}-c_{0j})(x_{mj^{\prime}}-c_{0j^{\prime}})=0$, for $j\neq j^{\prime}$; (iii) $\sum_{m=1}^{M}(x_{mj}-c_{0j})^{2}=Mr_{1}^{2}/3$.
Proof.
See Appendix C. ∎
Theorem 1 indicates that to achieve a minimum variance, the optimal measurement locations $\mathbf{x}_{m}$ should be distributed in an even and symmetric way about $\mathbf{c}_{0}$. One possible realization is to distribute the measurements uniformly on a ball with radius $Mr_{1}^{2}/3$ centered at $\mathbf{c}_{0}$.
The variance lower bound (18) consists of two terms where the first term $\sigma^{2}/M$ represents the error of the estimated $g(\mathbf{c}_{0})=\alpha$, and the second term represents the error of the estimated $\boldsymbol{\beta}$, i.e., the gradient $\nabla g(\mathbf{c}_{0})$ of the channel gain model (see Appendix D). Both terms decrease as the number of measurements $M$ increases. In addition, the error of $\boldsymbol{\beta}$ decreases as the measurement radius $r_{1}$ increases.
To derive the optimal measurement range $r_{1}$, we analyze the MSE of the locally constructed channel $\hat{g}(\mathbf{x})$ as follows.
Theorem 2 (MSE of the estimated channel gain).
Suppose that $\{\mathbf{x}_{m}\}$ for $m=1,2,\dots,M$ satisfy conditions (i)-(iii) in Theorem 1. For location $\mathbf{x}$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$, the MSE of the locally constructed channel $\hat{g}(\mathbf{x})$ is upper bounded as
$\mathbb{E}\left\{\left(\hat{g}(\mathbf{x})-g(\mathbf{x})\right)^{2}\right\}\leq$ $\frac{\sigma^{2}}{M}\left(1+\frac{3r_{0}^{2}}{r_{1}^{2}}\right)$ (19)
$+\frac{L_{g}^{2}}{4}\left(r_{1}^{2}+3r_{0}r_{1}+r_{0}^{2}\right)^{2}.$
In addition, if $\mathbf{x}_{m}$ also satisfies $\|\mathbf{x}_{m}-\mathbf{c}_{0}\|^{2}=r_{1}^{2}$, and if $g(\mathbf{x})$ can be locally approximated by a second order model, [^2] Under free-space propagation, the parameter $L^{\prime}_{g}$ in Theorem 2 can be calculated as $3.5\times 10^{-3}$ at a propagation distance of $50$ meters. i.e., $g(\mathbf{x})\approx g(\mathbf{c}_{0})+\bm{\beta}^{T}(\mathbf{x}-\mathbf{c}_{0})+L^{\prime}_{g}\|\mathbf{x}-\mathbf{c}_{0}\|^{2}/2$, then the MSE is approximately by
$\mathbb{E}\left\{\left(\hat{g}(\mathbf{x})-g(\mathbf{x})\right)^{2}\right\}$ $\approx\frac{\sigma^{2}}{M}\left(1+\frac{3r_{0}^{2}}{r_{1}^{2}}\right)+\frac{(L_{g}^{\prime})^{2}}{4}\left(r_{1}^{2}+r_{0}^{2}\right)^{2}.$ (20)
Proof.
See Appendix D. ∎
The result in Theorem 2 demonstrates a trade-off in the measurement range $r_{1}$. When the channel model $g(\mathbf{x})$ has a small Lipschitz constant $L_{g}$ in (2) and (3), corresponding to a small curvature, a large range $r_{1}$ is preferred for a small variance in (19), leading to a small MSE. When the channel model $g(\mathbf{x})$ has a large $L_{g}$, a small range $r_{1}$ is preferred, because the linear model (16) becomes less accurate in the range $r_{0}$ for a large $L_{g}$.
The MSE upper bound in Theorem 2 implies an optimal choice of the measurement range $r_{1}$. By minimizing the approximated MSE (20), Table I shows some numerical examples on the optimal choice of $r_{1}$ under different values of $M$ and $r_{0}$, where $\sigma=5$ dB and the parameter $L_{g}^{\prime}$ is obtained from a free-space propagation model evaluated for a neighborhood at a propagation distance of $50$ meters.
Table I: Optimal choice of $r_{1}$ [meter] under different $r_{0}$ [meter] and $M$
$r_{1}^{*}$ $M=40$ $M=60$ $M=80$ $M=100$
$r_{0}=10$ $17$ $16$ $15$ $14$
$r_{0}=20$ $20$ $18$ $17$ $16$
$r_{0}=30$ $21$ $20$ $18$ $17$
III-E Measurement Trajectory Design
Here, we construct the measurement trajectory $\mathbf{x}_{\text{r}}(t)$ to meet conditions (i)–(iii) in Theorem 1 for achieving a small error in locally constructing $\bm{g}(\mathbf{x})$ along the search $\mathbf{x}_{\text{s}}(t)$.
For the ease of elaboration, consider that the exploration direction is given by $\dot{\mathbf{x}}_{\text{s}}=\mathbf{s}=[0,1,0]^{\text{T}}$ for a piece of trajectory $\mathbf{x}_{\text{s}}(t)$ centered at $\mathbf{c}_{0}=[0,0,0]^{\text{T}}$ as illustrated in Fig. 2. One can construct a horizontal cylinder with length $2r_{1}/\sqrt{3}$ and radius $\sqrt{2/3}r_{1}$, where the measurement range $r_{1}$ is chosen according to Theorem 2 for a good construction performance within a one-step exploration range $r_{0}$ obtained while solving the equipotential surface tracking problem (13). The orientation of the cylinder is given by $\mathbf{s}$. It follows that if one uniformly samples along the circumferences of the top and bottom bases of the cylinder as shown by the blue dots in Fig. 2, the resulting sampling locations $\{\mathbf{x}_{m}\}$ satisfy all conditions in Theorem 1.
Note that the above trajectory only visits two distinct positions in the direction $\mathbf{s}$. Alternatively, along the search trajectory $\mathbf{x}_{\text{s}}(t)=[0,vr_{1}(t-M/2),0]^{\text{T}}$ for a speed parameter $v$, one may consider an alternating spiral trajectory $\mathbf{x}(t)=\mathbf{x}_{\text{s}}(t)+\mathbf{x}_{\text{r}}(t)$ with $\mathbf{x}_{\text{r}}(t)=[x_{\text{r}1}(t),0,x_{\text{r}3}(t)]^{\text{T}}$, where
$\begin{cases}x_{\text{r}1}(t)=\sqrt{2/3}r_{1}\cos(\omega t)\\ x_{\text{r}3}(t)=\sqrt{2/3}r_{1}\sin(\omega t)(-1)^{\lfloor\omega t/(2\pi)\rfloor}\end{cases}$ (21)
where $\omega=4k\pi/M$ and $v=2/\sqrt{M^{2}-1}$, with $k$ being a natural number, typically $k=1$ (see Fig. 2 (a)). One can easily verify that if we sample at $t=m-1/2$, i.e., $\mathbf{x}_{m}=\mathbf{x}(m-1/2)$, for $m=1,2,\dots,M$, then conditions (i)–(iii) in Theorem 1 are satisfied. As a result, the MSE of the locally reconstructed channel $\hat{g}_{k}(\mathbf{x})$ is upper bound by (19).
Figure 2: (a) An alternating spiral trajectory (orange dots) that satisfies conditions (i)–(iii) in Theorem 1. (b) A spiral trajectory that satisfies conditions (i) and (iii) in Theorem 1 and it is smooth.
IV Trajectory Design for Searching Optimal LOS Position on the Equipotential Surface
When the search is constrained on the equipotential surface $\mathcal{S}$ where it holds that $F(\bm{g}(\mathbf{x}))=f_{0}(g_{0}(\mathbf{x}))-F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p}^{*}(\mathbf{x}))=0$, the original problem $\mathscr{P}$ becomes
$\mathop{\mbox{maximize}}\limits_{\mathbf{x},\mathbf{p}}$ $\quad F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ (22)
[…]
$\quad H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0,n=1,2,\dots,N.$
It is very challenging to handle the constraint $\mathbf{x}\in\mathcal{S}$, especially when the analytical form of $\mathcal{S}$ is not available. Some classical approaches may consider projection-type algorithms, where the position $\mathbf{x}(t)$ is projected back to $\mathcal{S}$ whenever $\mathbf{x}(t)$ is off $\mathcal{S}$, e.g., via the trajectory developed in Section III. However, such a projection-type search is not suitable for UAV trajectory design except for an initialization phase, because frequent projections may cost a large amount of maneuver energy for the UAV. Therefore, it is desired that the UAV only moves on $\mathcal{S}$.
In this section, we develop a search trajectory $\mathbf{x}_{\text{s}}(t)$ sticking on the equipotential plane $\mathcal{S}$ without projections. The challenge is that a perturbation on $\mathbf{x}_{\text{s}}(t)$ may change the channel gain $g_{k}(\mathbf{x}_{\text{s}})$, and hence, the optimal resource allocation $\mathbf{p}^{*}(\mathbf{x}_{\text{s}})$, possibly resulting in $F(\bm{g}(\mathbf{x}_{\text{s}}))\neq 0$. We tackle this challenge via the perturbation theory.
IV-A Trajectory on the Equipotential Surface
[…]
For the problem (22), denote the Lagrangian function as
$L(\mathbf{p},\boldsymbol{\lambda};\bm{g}_{\text{u}}(\mathbf{x}))=F_{\text{u}}(\mathbf{p};\bm{g}_{\text{u}}(\mathbf{x}))-{\textstyle\sum_{n=1}^{N}}\lambda_{n}H_{n}(\mathbf{p};\bm{g}_{\text{u}}(\mathbf{x}))$
where $\boldsymbol{\lambda}=[\lambda_{1},\lambda_{2},\dots,\lambda_{N}]^{\text{T}}$. The Karush-Kuhn-Tucker (KKT) conditions are written as
$\mathbf{J}(\mathbf{p}(\mathbf{x}),\bm{\lambda}(\mathbf{x});\bm{g}_{\text{u}}(\mathbf{x}))\triangleq\left[\begin{array}[]{c}\nabla_{\mathbf{p}}L(\mathbf{p},\boldsymbol{\lambda};\bm{g}_{\text{u}}(\mathbf{x}))\\ \lambda_{1}H_{1}(\mathbf{p};\bm{g}_{\text{u}}(\mathbf{x}))\\ \lambda_{2}H_{2}(\mathbf{p};\bm{g}_{\text{u}}(\mathbf{x}))\\ \vdots\\ \lambda_{N}H_{N}(\mathbf{p};\bm{g}_{\text{u}}(\mathbf{x}))\end{array}\right]=0$ (23)
together with $\lambda_{n}\geq 0$ and $H_{n}(\mathbf{p}(\mathbf{x});\bm{g}_{\text{u}}(\mathbf{x}))\leq 0$ for all $n$. It is known that for a strictly convex problem, there is a unique solution $\{\mathbf{p}^{*},\boldsymbol{\lambda}^{*}\}$ to $\mathbf{J}(\mathbf{p}(\mathbf{x}),\bm{\lambda}(\mathbf{x});\bm{g}_{\text{u}}(\mathbf{x}))=0$ while satisfying $\lambda_{n}\geq 0$ and $H_{n}(\mathbf{p}(\mathbf{x});\bm{g}_{\text{u}}(\mathbf{x}))\leq 0$.
Since (23) and $F(\bm{g}(\mathbf{x}))=0$ are expected to be satisfied for all $\mathbf{x}(t)$, $t\geq 0$, we must have
$\begin{cases}\frac{\text{d}}{\text{d}t}\mathbf{J}(\mathbf{p}(\mathbf{x}(t)),\bm{\lambda}(\mathbf{x}(t));\bm{g}_{\text{u}}(\mathbf{x}(t)))&=0\\ \frac{\text{d}}{\text{d}t}F(\bm{g}(\mathbf{x}(t)))&=0\end{cases}$ (24)
which leads to
$\begin{cases}\nabla_{\mathbf{p}}\mathbf{J}^{\text{T}}\dot{\mathbf{p}}^{*}+\nabla_{\boldsymbol{\lambda}}\mathbf{J}^{\text{T}}\dot{\bm{\lambda}}^{*}+\nabla_{\bm{g}_{\text{u}}}\mathbf{J}^{\text{T}}\nabla\bm{g}_{\text{u}}^{\text{T}}\dot{\mathbf{x}}&=0\\ \nabla_{\mathbf{p}}F^{\text{T}}\dot{\mathbf{p}}^{*}+\nabla_{\bm{g}}F^{\text{T}}\nabla\bm{g}^{\text{T}}\dot{\mathbf{x}}&=0\end{cases}$ (25)
where we use the notation $\dot{\mathbf{p}}^{*}=\text{d}\mathbf{p}^{*}(t)/\text{d}t$, $\dot{\bm{\lambda}}^{*}=\text{d}\bm{\lambda}^{*}(t)/\text{d}t$, and $\dot{\mathbf{x}}=\text{d}\mathbf{x}(t)/\text{d}t$.
The above dynamical system (25) specifies a motion $\dot{\mathbf{x}}$ on the equipotential surface $\mathcal{S}$ with two spatial degrees of freedom. Suppose that the search is further constrained on a plane that intersects with $\mathcal{S}$. Denote $\mathbf{q}$ as the normal vector of the search plane, i.e., $\mathbf{q}^{\text{T}}\dot{\mathbf{x}}=0$. Then, the dynamic of the trajectory satisfies
$\left[\begin{array}[]{ccc}\nabla_{\mathbf{p}}\mathbf{J}^{\text{T}}&\nabla_{\boldsymbol{\lambda}}\mathbf{J}^{\text{T}}&\nabla_{\bm{g}_{\text{u}}}\mathbf{J}^{\text{T}}\nabla\bm{g}_{\text{u}}^{\text{T}}\\ \nabla_{\mathbf{p}}F^{\text{T}}&\mathbf{0}&\nabla_{\bm{g}}F^{\text{T}}\nabla\bm{g}^{\text{T}}\\ \mathbf{0}&\mathbf{0}&\mathbf{q}^{\text{T}}\\ \mathbf{0}&\mathbf{0}&\mathbf{v}^{\text{T}}\end{array}\right]\left[\begin{array}[]{c}\dot{\mathbf{p}}^{*}\\ \dot{\bm{\lambda}}^{*}\\ \dot{\mathbf{x}}\end{array}\right]$ $=\left[\begin{array}[]{c}\mathbf{0}\\ 1\end{array}\right]$ (35)
where the vector $\mathbf{v}$ can be randomly chosen and the last equality $\mathbf{v}^{\text{T}}\dot{\mathbf{x}}=1$ is to avoid the trivial solution $\dot{\mathbf{x}}=\mathbf{0}$, [^3] The physical meaning of $\mathbf{v}$ is to control the speed $\dot{\mathbf{x}}$ projected on the direction $\mathbf{v}$. For example, $\mathbf{v}=[0,0,1]^{\text{T}}$ specifies the speed $\dot{x}_{3}=1$. and therefore, the system of equations (35) is completely determined.
To derive a closed-form expression for $\dot{\mathbf{x}}$, let $\mathbf{A}_{1}=[\begin{array}[]{cc}\nabla_{\mathbf{p}}\mathbf{J}^{\text{T}}&\nabla_{\boldsymbol{\lambda}}\mathbf{J}^{\text{T}}\end{array}]$, $\mathbf{A}_{2}=\nabla_{\bm{g}_{\text{u}}}\mathbf{J}^{\text{T}}\nabla\bm{g}_{\text{u}}^{\text{T}}$,
$\mathbf{A}_{3}=\left[\begin{array}[]{cc}\nabla_{\mathbf{p}}F^{\text{T}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right],\quad\mathbf{A}_{4}(\mathbf{q})=\left[\begin{array}[]{c}\nabla F^{\text{T}}\nabla\bm{g}^{\text{T}}\\ \mathbf{q}^{\text{T}}\\ \mathbf{v}^{\text{T}}\end{array}\right]$ (36)
and $\mathbf{e}_{3}=[0,0,1]^{\text{T}}$. Using the block matrix inversion lemma, the dynamic $\dot{\mathbf{x}}$ as the solution to (35) can be derived as
$\dot{\mathbf{x}}=\mathcal{A}(\mathbf{x};\mathbf{q})\triangleq\left(\mathbf{A}_{4}(\mathbf{q})-\mathbf{A}_{3}\mathbf{A}_{1}^{-1}\mathbf{A}_{2}\right)^{-1}\mathbf{e}_{3}.$ (37)
Figure 3: (a) LOS discovery trajectory on the equipotential surface. (b) Search directions in Phase $1$ and $2$.
IV-B LOS Discovery on the Equipotential Surface
The dynamical system (35) requires specifying a search direction determined by the normal vector $\mathbf{q}$ in (35). We present a search strategy to adaptively determine the search direction via the normal vector $\mathbf{q}$ to discover the best LOS opportunity that solves problem (22). The strategy is based on the following two properties.
First, consider the curve $f_{0}(g_{0}(\mathbf{x}))=C$ on the equipotential surface $\mathbf{x}\in\mathcal{S}$, as illustrated by the dashed curve in Fig. 3(a). It follows that the lower the curve, i.e., with a smaller radius, the larger the objective value $f_{0}(g_{0}(\mathbf{x}))=F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p}^{*}(\mathbf{x}))$, because $\mathbf{x}$ is closer to the BS resulting in a larger channel gain $g_{0}(\mathbf{x})$. Second, there is an upward invariant property for the full LOS region $\tilde{\mathcal{D}}$ due to the environment model in Section II-A, i.e., if $\mathbf{x}\notin\tilde{\mathcal{D}}$, the locations below $\mathbf{x}$ are also in NLOS.
These observations inspire the following search strategy for an LOS discovery trajectory $\mathbf{x}_{\text{s}}(t)$.
  • Phase 1: if $\mathbf{x}_{\text{s}}(t)\in\mathcal{S}\cap\tilde{\mathcal{D}}$, then one should decrease the altitude of $\mathbf{x}_{\text{s}}(t)$ to discover a larger $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))$. Thus, the search direction $\dot{\mathbf{x}}$ should lie on the tangent plane of the equipotential surface $\mathcal{S}$ and be as close to the downward vector $-\mathbf{e}_{3}=[0,0,-1]^{\text{T}}$ as possible. Analytically, the normal vector $\mathbf{q}$ in (35) of the search plan that contains $\dot{\mathbf{x}}$ should be orthogonal to both the normal vector $\text{grad}\,F(\bm{g}(\mathbf{x}_{\text{s}}))=\nabla F(\bm{g}(\mathbf{c}_{0}))^{\text{T}}\mathbf{G}(\mathbf{c}_{0})$ for $\mathcal{S}$ and the downward vector $-\mathbf{e}_{3}$, i.e.,
    $\mathbf{q}_{1}(\mathbf{x}_{\text{s}})=\frac{\text{grad}\,F(\bm{g}(\mathbf{x}_{\text{s}}))\times(-\mathbf{e}_{3})}{\|\text{grad}\,F(\bm{g}(\mathbf{x}_{\text{s}}))\times(-\mathbf{e}_{3})\|_{2}}$ (38)
    where $\mathbf{a}\times\mathbf{b}$ denotes the cross product of $\mathbf{a}$ and $\mathbf{b}$.
  • Phase 2: if $\mathbf{x}_{\text{s}}(t)\in\mathcal{S}$ but $\mathbf{x}\notin\tilde{\mathcal{D}}$, one explores the equipotential surface following the curve $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=C$ to discover an LOS opportunity. Analytically, the curve $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=C$ satisfies the following ODE $\frac{\text{d}}{\text{d}t}f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=\nabla f_{0}\nabla g_{0}^{\text{T}}\dot{\mathbf{x}}_{\text{s}}=0$, and thus, the normal vector $\mathbf{q}$ is given by
    $\mathbf{q}_{2}(\mathbf{x}_{\text{s}})=\text{$\frac{\nabla f_{0}(g_{0}(\mathbf{x}_{\text{s}}))\nabla g_{0}(\mathbf{x}_{\text{s}})^{\text{T}}}{\|\nabla f_{0}(g_{0}(\mathbf{x}_{\text{s}}))\nabla g_{0}(\mathbf{x}_{\text{s}})^{\text{T}}\|_{2}}$}.$ (39)
IV-C Superposed Trajectory via ODEs
IV-C1 The Superposed Trajectory
Here, we combine the search trajectory $\mathbf{x}_{\text{s}}(t)$ with the measurement trajectory $\mathbf{x}_{\text{r}}(t)$ developed in Section III-E, assuming the initial state satisfies $\mathbf{x}_{\text{s}}(0)\in\mathcal{S}$.
First, from (15) and (37), the combined search trajectory is given by $\dot{\mathbf{x}}_{\text{s}}=\mathcal{A}(\mathbf{x}_{\text{s}}(t);\mathbf{q}(\mathbf{x}_{\text{s}}(t)))+\mu_{\text{v}}\mathcal{V}(\mathbf{x}_{\text{s}}(t))$, where the first term is to search on the equipotential surface $\mathcal{S}$ according to the two exploration phases (38) and (39), and the second term is to track the $\mathcal{S}$ in case $\mathbf{x}_{\text{s}}(t)$ deviates from it due to implementation issues.
Second, consider the measurement trajectory $\mathbf{x}_{\text{r}}(t)=r[\cos(\omega t),\,0,\,\sin(\omega t)]^{\text{T}}$ developed in Section III-E, which forms a circle on the plane with a normal vector $\mathbf{e}_{2}=[0,1,0]^{\text{T}}$. Then, given the search direction $\dot{\mathbf{x}}_{\text{s}}$, one can construct a rotation matrix $\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})$ that rotates the coordinate system with the reference direction $\mathbf{e}_{2}$ to a new coordinate system with the reference direction $\dot{\mathbf{x}}_{\text{s}}/\|\dot{\mathbf{x}}_{\text{s}}\|_{2}$. The rotation matrix $\mathbf{R}(\mathbf{s})$ to the reference direction $\mathbf{s}=[s_{1},s_{2},s_{3}]^{\text{T}}$ is found as
$\mathbf{R}(\mathbf{s})=\mathbf{I}-\frac{1}{\|\mathbf{s}\|_{2}}\left[\begin{array}[]{ccc}\frac{s_{1}^{2}}{\|\mathbf{s}\|_{2}+s_{2}}&s_{1}&\frac{-s_{1}s_{3}}{\|\mathbf{s}\|_{2}+s_{2}}\\ -s_{1}&\|\mathbf{s}\|_{2}-s_{2}&-s_{3}\\ \frac{-s_{1}s_{3}}{\|\mathbf{s}\|_{2}+s_{2}}&s_{3}&\frac{s_{3}^{2}}{\|\mathbf{s}\|_{2}+s_{2}}\end{array}\right].$
The dynamical equation for the superposed UAV search trajectory $\mathbf{x}(t)=\mathbf{x}_{\text{s}}(t)+\mathbf{x}_{\text{r}}(t)$ then becomes
$\dot{\mathbf{x}}$ $=\dot{\mathbf{x}}_{\text{s}}+\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})\dot{\mathbf{x}}_{\text{r}}+\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})\mathbf{x}_{\text{r}}(t)$ (40)
$\dot{\mathbf{x}}_{\text{s}}$ $=\mathcal{A}(\mathbf{x}_{\text{s}}(t);\mathbf{q}(\mathbf{x}_{\text{s}}(t)))+\mu_{\text{v}}\mathcal{V}(\mathbf{x}_{\text{s}}(t))$ (41)
where $\dot{\mathbf{x}}_{\text{r}}=\text{d}\mathbf{x}_{\text{r}}(t)/\text{d}t=r\omega[-\sin(\omega t),\,0,\,\cos(\omega t)]^{\text{T}}$ and $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})=\nabla\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})(\mathbf{I}_{3}\otimes\ddot{\mathbf{x}}_{\text{s}})$. Here, the operator $\nabla\mathbf{R}$ gives a matrix with $3\times 3$ blocks, where the $(i,j)$th block is $[\frac{\partial R_{ij}}{\partial x_{1}},\frac{\partial R_{ij}}{\partial x_{2}},\frac{\partial R_{ij}}{\partial x_{3}}]$, $\otimes$ is the Kronecker product, and $\ddot{\mathbf{x}}_{\text{s}}=\text{d}^{2}\mathbf{x}_{\text{s}}/\text{d}t^{2}$. In (40), the second term creates a spiral trajectory surrounding the main search route $\mathbf{x}_{\text{s}}(t)$ for collecting the channel measurement data, and the third term generates the adjustment due to the potential time variation of the search direction $\dot{\mathbf{x}}_{\text{s}}$.
IV-C2 Implementation
The analytical form of $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})$ in (40) requires the second-order derivative of the search trajectory $\mathbf{x}_{\text{s}}(t)$, which is not available. A simple solution is to use numerical approximations $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})\approx\frac{1}{\tau}(\mathbf{R}(\dot{\mathbf{x}}_{\text{s}}(t))-\mathbf{R}(\dot{\mathbf{x}}_{\text{s}}(t-\tau)))$ for a small enough $\tau>0$. Alternatively, we find the following approximations.
First, when the search $\mathbf{x}_{\text{s}}(t)$ remains in Phase 1 or Phase 2 as specified in Section IV-B, we likely have $\ddot{\mathbf{x}}_{\text{s}}\approx\mathbf{0}$, and thus the term $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})\mathbf{x}_{\text{r}}(t)$ can be simply ignored. This is because $\mathbf{x}_{\text{s}}(t)$ is a trajectory on the equipotential surface $\mathcal{S}$, and thus $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})$ represents the supplementary rotation due to the curvature of $\mathcal{S}$, which is relatively small in practical regime of interest compared to the other terms. [^4] For example, for a balancing problem where the equipotential plane $\mathcal{S}$ is a sphere under some mild conditions (Proposition 2), it is known that the curvature of a sphere with radius $R$ is $1/R^{2}$, which is small as $R$ is at the order of a hundred meters. Hence, the dynamical equation for $\mathbf{x}(t)$ in this case is approximated as
$\dot{\mathbf{x}}$ $\approx\dot{\mathbf{x}}_{\text{s}}+\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})\dot{\mathbf{x}}_{\text{r}}.$ (42)
Second, when the search $\mathbf{x}_{\text{s}}(t)$ needs to switch, for instance, from Phase 1 to Phase 2 at time $t=t_{1}$, the search direction needs to switch between $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})=\mathcal{A}(\mathbf{x}_{\text{s}}(t_{1}^{-});\mathbf{q}_{1}(\mathbf{x}_{\text{s}}(t_{1}^{-})))$ and $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})=\mathcal{A}(\mathbf{x}_{\text{s}}(t_{1}^{+});\mathbf{q}_{2}(\mathbf{x}_{\text{s}}(t_{1}^{+})))$, assuming $\mathcal{V}(\mathbf{x}_{\text{s}}(t))=0$ for simplicity. Thus, $\frac{\text{d}}{\text{d}t}\mathbf{R}(\dot{\mathbf{x}}_{\text{s}})$ does not exist as $\ddot{\mathbf{x}}_{\text{s}}$ does not exist, since $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})\neq\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})$. To circumvent this issue, a transition phase $t\in(t_{1},t_{1}+\tau)$ is needed, where without altering $\mathbf{x}_{\text{s}}(t)$, i.e., $\dot{\mathbf{x}}_{\text{s}}=0$, we smoothly switch $\dot{\mathbf{x}}_{\text{s}}$ from $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})$ to $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})$ using a linear transition
$\dot{\mathbf{x}}_{\text{t}}=\frac{\tau-t+t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})+\frac{t-t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+}),\quad t\in(t_{1},t_{1}+\tau)$ (43)
which yields $\ddot{\mathbf{x}}_{\text{t}}=\frac{1}{\tau}(\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})-\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-}))$. As a result, the dynamical equation for $\mathbf{x}(t)$ becomes, for $t\in(t_{1},t_{1}+\tau)$,
$\dot{\mathbf{x}}$ $=\mathbf{R}(\dot{\mathbf{x}}_{\text{t}})\dot{\mathbf{x}}_{\text{r}}+\frac{1}{\tau}\nabla\mathbf{R}(\dot{\mathbf{x}}_{\text{t}})(\mathbf{I}_{3}\otimes(\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})-\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})))\mathbf{x}_{\text{r}}(t).$ (44)
A sample implementation is summarized in Algorithm 1. As Algorithm 1 is an online search scheme, the following two metrics are essential for understanding its complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course. Specifically, given a common realization of channel models and a specified objective function for the balancing problem in Proposition 2, the trajectory length $L$ is found to be linear in some geometry parameters of the environment, i.e.,
$L\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}$ (45)
where $H_{0}$ is the initial search altitude, $R$ is the radius of $\mathcal{S}$ given in (11), and $r_{1}$ is the measurement range defined in Section III. Recall that $K$ is the number of users, $M$ is the number of measurements used for local channel model construction, and $N$ is the number of constraints for resource allocation. The per-step computational complexity is found as $O(KM+(K+N)^{3})$, which is numerically found to be handled in milliseconds. For detailed complexity derivations, see Appendix E: Supplementary Material 1.
Find a full-LOS initial position $\mathbf{x}_{0}\in\mathcal{S}\cap\tilde{\mathcal{D}}$. Denote $\mathbf{x}_{\text{r}}(t)=r[\cos(\omega t),\,0,\,\sin(\omega t)]^{\text{T}}$.
  1. 1.
    Initialization at $t=0$: $\tilde{\mathbf{x}}=\mathbf{x}_{0}$, $\mathbf{x}_{\text{s}}(0)=\mathbf{x}_{0}$, and $\mathbf{x}(0)=\mathbf{x}_{\text{s}}(0)+\mathbf{x}_{\text{r}}(0)$.
  2. 2.
    While $x_{3}(t)\geq H_{\text{min}}$:
  3. 3.
    Local channel construction: Collect channel measurements for each time slot $\triangle t$. Construct the local channel $\hat{g}_{k}(\mathbf{x})$ and obtain the estimate for $\nabla g_{k}(\mathbf{x})$ for each user based on model (16) with parameter (17) estimated from the past $M$ LOS measurements.
  4. 4.
    Update the UAV location $\mathbf{x}(t)$ according to the following cases:
    1. (a)
      Phase 1 (LOS): If $\mathbf{x}(t-\Delta t),\mathbf{x}(t)\in\tilde{\mathcal{D}}$,
      1. i.
        If $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))>f_{0}(g_{0}(\tilde{\mathbf{x}}))$, then update the optimal position $\tilde{\mathbf{x}}\leftarrow\mathbf{x}_{\text{s}}(t)$.
      2. ii.
        Compute the search direction $\mathbf{s}(t)=\mathcal{A}(\mathbf{x}_{\text{s}}(t);\mathbf{q}_{1}(\mathbf{x}_{\text{s}}(t)))+\mu_{\text{v}}\mathcal{V}(\mathbf{x}_{\text{s}}(t))$ from (14), (37), and (38).
      3. iii.
        Update $\mathbf{x}_{\text{s}}(t+\triangle t)=\mathbf{x}_{\text{s}}(t)+\mathbf{s}\triangle t$ and $\mathbf{x}(t+\triangle t)=\mathbf{x}(t)+(\mathbf{s}+\mathbf{R}(\mathbf{s})\dot{\mathbf{x}}_{\text{r}})\triangle t$ according to (40) and (42).
    2. (b)
      Phase 2 (NLOS): If $\mathbf{x}(t-\Delta t),\mathbf{x}(t)\neq\tilde{\mathcal{D}}$,
      1. i.
        Compute the search direction $\mathbf{s}(t)=\mathcal{A}(\mathbf{x}_{\text{s}}(t);\mathbf{q}_{2}(\mathbf{x}_{\text{s}}(t)))+\mu_{\text{v}}\mathcal{V}(\mathbf{x}_{\text{s}}(t))$ from (14), (37), and (39).
      2. ii.
        The same as Step 4(a)iii.
    3. (c)
      Otherwise (Transition Phase):
      1. i.
        Let $t_{1}\leftarrow t-\triangle t$.
      2. ii.
        Compute $\dot{\mathbf{x}}_{\text{t}}$ and $\bm{\delta}=\dot{\mathbf{x}}$ respectively according to (43) and (44), by replacing “$\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})$” with $\mathbf{s}(t_{1})$ and replacing “$\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})$” with $\mathbf{s}(t_{1}+\triangle t)$.
      3. iii.
        Update $\mathbf{x}_{\text{s}}(t+\triangle t)=\mathbf{x}_{\text{s}}(t)$ and $\mathbf{x}(t+\triangle t)=\mathbf{x}(t)+\bm{\delta}\triangle t$.
      4. iv.
        Repeat from Step 4(c)ii until $t\geq t_{1}+\tau$.
  5. 5.
    End while; output $\tilde{\mathbf{x}}$ as the best position found on $\mathcal{S}\cap\tilde{\mathcal{D}}$.
Algorithm 1 Superposed LOS discovery and measurement collection trajectory with unknown user locations
V Numerical Results
In this section, we present the experimental findings conducted on two real 3D urban maps.
V-A Environment Setup and Scenarios
Two city maps of different areas in Beijing, China are used to evaluate the proposed scheme. As shown in Fig. 4, map A represents a sparse commercial area with the building coverage ratio (BCR) and floor area ratio (FAR) [27] as $18\%$ and $1.0$, respectively, while map B represents a dense residential area with the BCR and FAR as $33\%$ and $1.86$, respectively. The minimum flying altitudes $H_{\text{min}}$ are set as $29$ and $62$ meters on maps A and B, respectively, which correspond to the minimum height of the top $20\%$ tallest buildings. The users are placed uniformly at random in the non-building area for $2000$ repetitions, with the BS placed above an arbitrarily chosen building. For the proposed scheme, $M$, $\omega$, $r$, $\mu_{\text{v}}$, and $\tau$ are set as $100$, $\pi/25$, $25$, $1$, and $5$, respectively.
Two application scenarios are evaluated in our experiments. For a sum-rate application, one UAV is placed to establish LOS relay channels for $K$ ground users and a BS under decode-and-forward relaying. Consider the path loss model of millimeter wave cellular reported in [28, 29], where the deterministic LOS channel gain $g_{k}(\mathbf{x})$ is modeled as $g_{k}(\mathbf{x})=46.53+20.0\log_{10}d(\mathbf{x},\mathbf{u}_{k})$. The variance $\sigma^{2}$ of measurement uncertainty $\xi$ in (4) is set as $5$ dB. It is assumed that the mmW beam alignment has been done for every UAV position. For a balancing application, one UAV is placed to provide detection services for eight sensing targets and maintains a backhaul communication link with a BS simultaneously. Particularly, the objective functions for sensing are specified as SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=p_{k}g_{k}(\mathbf{x})/(\bar{N}_{0}W)$ where the noise power density $\bar{N}_{0}$ is set as $-150$ dBm/Hz, and the bandwidth $W$ is set as $1$ GHz.
Figure 4: Map A (left) is a sparse commercial area, and map B (right) is a dense residential area in Beijing, China.
The baseline schemes are listed below, and the exhaustive search schemes are implemented with a $5$-meter step size.
  • Exhaustive 3D search (Exh3D): This scheme performs an exhaustive search in 3D space above the area of interest.
  • Exhaustive 2D search (Exh2D) [30]: This scheme performs an exhaustive search at a height of $H_{\text{min}}+50$.
  • Statistical geometry (Statis) [31, 20]: The average channel gain from the UAV position $\mathbf{x}$ to the $k$th user is formulated as $\tilde{g}_{k}(\mathbf{x})=\text{P}_{\text{L}}(\mathbf{x})g_{k}(\mathbf{x})+(1-\text{P}_{\text{L}}(\mathbf{x}))(g_{k}(\mathbf{x})+\phi(\mathbf{x}))$ where the power penalty $\phi(\mathbf{x})$ for NLOS link is set as $-30$ dB, and $\text{P}_{\text{L}}(\mathbf{x})$ is the LOS probability at $\mathbf{x}$, which is defined as $\text{P}_{\text{L}}(\mathbf{x})=1/(1+a_{\text{e}}\times\text{exp}(-b_{\text{e}}(\text{arctan}(x_{3}/\sqrt{\|\mathbf{x}-\mathbf{u}_{k}\|_{2}^{2}-x_{3}^{2}})-a_{\text{e}})))$ where the parameters $a_{\text{e}}$ and $b_{\text{e}}$ are learned from maps.
  • Relaxed analytical geometry (RAG) [10]: This scheme models the city structure using polyhedrons and determines the LOS conditions via a set of constraints obtained from analytical geometry. The UAV position optimization problem is solved by Lagrangian relaxation and successive convex approximation (SCA).
Note that both the statistical geometry scheme and the relaxed analytical geometry scheme require user locations and channel model parameters, and hence, they serve for performance benchmarking only. Additionally, we test a genius-aided version of the proposed scheme with known user locations and channel models, and thus, the measurement range $r_{1}$ is set to 00. This is to set the benchmark for the best possible performance of searching on the equipotential surface.
V-B UAV-assisted Communication and Sensing
Figure 5: Normalized estimation error of channel gain $g(\mathbf{x})$ versus measurement range $r_{1}$.
Figure 6: Normalized estimation error of local gradient of channel gain, i.e., $\nabla g(\mathbf{x})$, versus measurement range $r_{1}$.
Fig. 5 shows the normalized estimation error of channel gain $g(\mathbf{x})$ under different settings of $M$ and $r_{1}$. It is found that the error in estimating $g(\mathbf{x})$ exhibits a monotonic decrease w.r.t. the increase of $M$ while it is not monotonic w.r.t. $r_{1}$. This observation suggests the existence of an optimal value for $r_{1}$, e.g., $r_{1}\approx 18$ when $M=60$, which confirms the analytical results in Theorem 2. Fig. 6 demonstrates the normalized estimation error of the gradient of channel gain $\nabla g(\mathbf{x})$. It is observed that a small measurement range $r_{1}$ leads to a large estimation error on the local gradient of $g(\mathbf{x})$ as the measurement is not geographically diverse. Yet, when $r_{1}$ is too large, the first-order local LOS model $\hat{g}(\mathbf{x})$ becomes less accurate, leading to a slight increase of the estimation error.
Table II: Capacity on two maps for a sum-rate application [${\rm Gbps}$]
Statis Exh2D RAG Proposed Genius-aided Exh3D
Map A 3.21 3.52 3.72 3.82 3.84 3.85
Map B 1.70 1.84 2.08 2.17 2.24 2.30
Figure 7: System capacity versus transmit power of UAV
Table II summarizes the system capacity, i.e., $\min\{f_{0}(g_{0}(\mathbf{x})),\sum_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k}\text{)}\}\}$, of different schemes on two maps. The proposed scheme achieves over $94\%$ to that of the exhaustive 3D search, only 3% below the genius-aided scheme which requires perfect user locations and channel parameters. Such a result suggests that the globally optimal solution likely resides on the equipotential surface, and the proposed scheme can search near the locally reconstructed equipotential surface without significant deviation. Although the relaxed analytical geometry scheme achieves over $90\%$ to the exhaustive 3D search over two maps, it requires complete knowledge of city topology and additional computational cost for polygonal approximation of buildings. The suboptimal performance of the exhaustive 2D search is primarily due to its inability to leverage height flexibility. When compared to other schemes with LOS guarantee, the poor performance of the statistical geometry scheme emerges from the inherent uncertainty of the LOS status of potentially best UAV positions found through this scheme.
Fig. 7 demonstrates the system capacity under varying UAV transmit powers, with the transmit power of BS set to $K$ times that of the UAV. It is observed that the proposed scheme closely approaches the performance of the exhaustive 3D search regardless of transmit power. Although a denser map makes it harder to find the globally optimal solution, the proposed scheme still achieves over $94\%$ to the exhaustive 3D search on map B while it achieves over $99\%$ to the exhaustive 3D search on map A under $P_{\text{T}}=30$ dBm.
Fig. 8 shows the system capacity under different user numbers in the sum-rate application. It is found that the proposed scheme achieves over $98\%$ to the exhaustive 3D search on map A under all tested user numbers, while it achieves up $90\%$ on map B, which represents a denser urban area. The reason is that denser urban topology and more users result in much more complicated LOS distributions. Thus, gathering a sufficient number of LOS measurements in a local area for channel estimation becomes more challenging, resulting in increased construction error of the equipotential surface.
Figure 8: System capacity versus number of users
Figure 9: CDF versus $\text{SNR}=\min_{k\in\mathcal{K}}\{p_{k}^{*}g_{k}(\mathbf{x})/N_{0}\}$ in a UAV-assisted sensing and communication application.
Table III: Comparison of Average Trajectory Length [kilometer]
Genius-aided Proposed Exh2D Exh3D
Map A 0.2604 3.272 1920 42240
Map B 0.2428 3.051 1920 42240
In Fig. 9, we examine the proposed scheme in a balancing application. As observed, the proposed scheme substantially increases the SNR of the weakest sensing link over the statistical geometry scheme. In particular, it is observed that the cumulative distribution function (CDF) curve of the proposed scheme closely coincides with that of the exhaustive 3D search on map B, numerically confirming that the proposed scheme can find the near-optimal solution in complex urban environments.
Table III summarizes the average trajectory lengths required by the four online schemes on two maps. Notably, the genius-aided scheme with known user locations and channel models requires only a few hundred meters to approximate a near-optimal solution, confirming the efficiency of the LOS discovery strategy proposed in Section IV-B. The proposed scheme demands a longer trajectory of approximately $3$ kilometers due to its reliance on spiral trajectories for data collection during the search process. Despite this, the proposed scheme still demonstrates considerable efficacy, significantly reducing search complexity compared to the exhaustive search schemes.
VI Conclusion
In this paper, we developed an efficient online trajectory for UAV placement without prior knowledge of user locations, channel model parameters, or terrain structure. We analytically characterized the equipotential surface and proposed an LOS discovery strategy on it, utilizing perturbation theory to guide the UAV search direction. Furthermore, we introduced spiral trajectories to construct local channel models in the LOS regime using local polynomial regression, independent of user locations. After deriving the optimal measurement pattern, we minimized the MSE of the locally estimated channel gain and determined the optimal measurement range. Experimental results on real urban maps demonstrated that our approach achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search, offering a practical solution for UAV-assisted sensing and communication in complex environments.
Appendix A Proof of Proposition 1
First, the optimal power allocation in the balancing problem is expressed as
$p_{k}^{*}=\frac{P_{\text{T}}}{g_{k}(\mathbf{x})\sum_{k^{\prime}\in\mathcal{K}}(1/g_{k^{\prime}}(\mathbf{x}))}.$ (46)
which is achieved when all of the users share the same received SNR, i.e., $p_{1}g_{1}(\mathbf{x})=p_{2}g_{2}(\mathbf{x})=\cdots=p_{K}g_{K}(\mathbf{x})$. Otherwise, the UAV can enhance the worst link performance by allocating more power to that link. Since there is a maximum power constraint $\sum_{k=1}^{K}p_{k}\leq P_{\text{T}}$, the optimal solution $p_{k}^{*}$ in (46) can be derived. Second, the condition in (10) can be equivalently rewritten as $F(\bm{g}(\mathbf{x}_{0}^{\text{m}}))F(\bm{g}(\mathbf{x}_{\text{\text{u}}}^{\text{m}}))\leq 0$. Third, according to the Bolzano-Cauchy theorem and the continuity of $F(\bm{g}(\mathbf{x}))$, there exists an equipotential point $\mathbf{x}$ satisfying $F(\bm{g}(\mathbf{x}))=0$, confirming the existence of the equipotential surface.
Appendix B Proof of Proposition 2
Recall that the optimal power allocation in the balancing problem is derived in (46). Thus, the objective function of the UAV-user link is rewritten as
$F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p}^{*})=\log_{2}(1+P_{\text{T}}/{\textstyle\sum_{k^{\prime}\in\mathcal{K}}}(1/g_{k^{\prime}}(\mathbf{x}))).$ (47)
Given $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$, the equation $F(\bm{g}(\mathbf{x}))=0$ is simplified as $P_{\text{T}}d^{2}(\mathbf{x},\mathbf{u}_{0})-{\textstyle P_{0}\sum_{k^{\prime}\in\mathcal{K}}d^{2}(\mathbf{x},\mathbf{u}_{k})}=0$ which can be rewritten as a spherical equation with specified center and radius in (11).
Appendix C Proof of Theorem 1
We first derive the following lemma on trace of the inverse of a positive definite and symmetric matrix.
Lemma 1 (Trace of the inverse of a positive definite and symmetric matrix).
Let $\mathbf{A}\in\mathcal{M}_{N}$ be a positive definite symmetric matrix, and $a_{ii}$ be the $i$th diagonal element. Then,
$\text{{\rm tr}}\{\mathbf{A}^{-1}\}\geq\sum_{i=1}^{N}a_{ii}^{-1}$
with the equality achieved for $a_{ij}=0$, $\forall i\neq j$.
Proof.
See Appendix F: Supplementary Material 2. ∎
According to [25, 26], the variance of $\boldsymbol{\theta}^{\text{(e)}}$ is given by $\text{tr}\{\sigma^{2}(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$. Using the results in Lemma 1, one can minimize $\text{tr}\{(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$ with the optimal conditions on $\tilde{\boldsymbol{\mathbf{X}}}$.
Since $\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}$ is invertible, none of its eigenvalues is equal to 00. Given the definition of $\tilde{\mathbf{X}}$, all the primary minors of $\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}$ are no less than 00. Hence, $\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}$ is a positive definite symmetric matrix. According to Lemma 1,
$\text{tr}\{(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$ $\geq\frac{1}{M}+\sum_{j=1}^{3}\frac{1}{\sum_{m=1}^{M}(x_{mj}-c_{0j})^{2}}$
with the minimum achieved when $\sum_{m=1}^{M}(x_{mj}-c_{0j})=0$, for $\forall j\in\{1,2,3\}$, and $\sum_{m=1}^{M}(x_{mj}-c_{0j})(x_{mj^{\prime}}-c_{0j^{\prime}})=0$, for $\forall j,j^{\prime}\in\{1,2,3\}$ and $j\neq j^{\prime}$.
Using the Arithmetic Geometric Mean Inequality (AGMI) and the fact that $\sum_{j=1}^{3}\sum_{m=1}^{M}(x_{mj}-c_{0j})^{2}\leq Mr_{1}^{2}$, we have
$\text{tr}\{(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$ $\frac{1}{M}+\frac{3}{\sqrt[3]{\prod_{j=1}^{3}\sum_{m=1}^{M}(x_{mj}-c_{0j})^{2}}}$
$=\frac{1}{M}+\frac{9}{Mr_{1}^{2}}$
with the minimum achieved when $\sum_{m=1}^{M}(x_{mj}-c_{0j})^{2}=Mr_{1}^{2}/3$ for $j\in\{1,2,3\}$. Thus, the lower bound of $\text{tr}\{\sigma^{2}(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$ is correspondingly obtained.
Appendix D Proof of Theorem 2
The MSE of the estimated channel gain at $\mathbf{x}$ is given by
$\mathbb{E}\left\{\left(\hat{g}(\mathbf{x})-g(\mathbf{x})\right)^{2}\right\}$ $=\mathbb{V}\left\{\hat{g}(\mathbf{x})\right\}+\left(\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\right)^{2}$ (48)
Next, the bounds of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ and $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$ are derived as below.
1) Derivation of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$: According to the proof of Theorem 1, the variances of $\hat{\alpha}$ and $\hat{\boldsymbol{\beta}}$ are given by $\mathbb{V}\{\hat{\alpha}\}=\sigma^{2}/M$, and $\mathbb{V}\{\hat{\beta}_{j}\}=3\sigma^{2}/Mr_{1}^{2}$. Thus, $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ is given by
$\mathbb{V}\left\{\hat{g}(\mathbf{x})\right\}$ $=\mathbb{V}\left\{\hat{\alpha}+\hat{\boldsymbol{\beta}}^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})\right\}=\frac{\sigma^{2}}{M}+\frac{3\sigma^{2}r_{0}^{2}}{Mr_{1}^{2}}.$ (49)
2) Derivation of $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$: Since $\{\mathbf{x}_{m}\}$ satisfies conditions (i)-(iii) in Theorem 1, $(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}$ is simplified as
$(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}=\frac{1}{M}\left[\begin{array}[]{cc}1&\mathbf{0}\\ \mathbf{0}&\frac{3}{r_{1}^{2}}\mathbf{I}\end{array}\right]$
where $\mathbf{I}$ is a $3\times 3$ unit matrix. Therefore, $\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}$ is derived as
$\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}=\frac{1}{M}\left[\begin{array}[]{ccc}1&\cdots&1\\ \frac{3}{r_{1}^{2}}(\mathbf{x}_{1}-\mathbf{c}_{0})&\cdots&\frac{3}{r_{1}^{2}}(\mathbf{x}_{M}-\mathbf{c}_{0})\end{array}\right].$
Recall that $g(\mathbf{x})$ satisfies that Lipschitz condition in (2) and (3). Hence, we have
$y_{m}$ $=g(\mathbf{x}_{m})+\xi_{m}$
$=\left[\begin{array}[]{cc}1&(\mathbf{x}_{m}-\mathbf{c}_{0})^{\text{T}}\end{array}\right]\left[\begin{array}[]{c}g(\mathbf{c}_{0})\\ \nabla g(\mathbf{c}_{0})\end{array}\right]+e(\mathbf{x}_{m},\mathbf{c}_{0})+\xi_{m}$
where $|e(\mathbf{x}_{m},\mathbf{c}_{0})|\leq L_{g}d^{2}(\mathbf{x}_{m},\mathbf{c}_{0})/2\leq L_{g}r_{1}^{2}/2$, and $\xi_{m}\sim N(0,\sigma^{2})$ is the measurement noise at $\mathbf{x}_{m}$.
Recall that $\mathbf{y}=[y_{1},y_{2},\dots,y_{M}]^{\text{T}}$. Then, we have
$\mathbf{y}$ $=\left[\begin{array}[]{cc}1&(\mathbf{x}_{1}-\mathbf{c}_{0})^{\text{T}}\\ 1&(\mathbf{x}_{2}-\mathbf{c}_{0})^{\text{T}}\\ \vdots&\vdots\\ 1&(\mathbf{x}_{M}-\mathbf{c}_{0})^{\text{T}}\end{array}\right]\left[\begin{array}[]{c}g(\mathbf{c}_{0})\\ \nabla g(\mathbf{c}_{0})\end{array}\right]+\left[\begin{array}[]{c}e(\mathbf{x}_{1},\mathbf{c}_{0})+\xi_{1}\\ e(\mathbf{x}_{2},\mathbf{c}_{0})+\xi_{2}\\ \vdots\\ e(\mathbf{x}_{M},\mathbf{c}_{0})+\xi_{M}\end{array}\right]$
$=\tilde{\boldsymbol{\mathbf{X}}}\left[\begin{array}[]{c}g(\mathbf{c}_{0})\\ \nabla g(\mathbf{c}_{0})\end{array}\right]+\mathbf{e}+\boldsymbol{\xi}$
where $\mathbf{e}=[e(\mathbf{x}_{1},\mathbf{c}_{0}),e(\mathbf{x}_{2},\mathbf{c}_{0}),\dots,e(\mathbf{x}_{M},\mathbf{c}_{0})]^{\text{T}}$, and $\bm{\xi}=[\xi_{1},\xi_{2},\dots,\xi_{M}]^{\text{T}}$ satisfying $\mathbb{E}\{\bm{\xi}\}=\mathbf{0}$.
Denote $\bar{\bm{\theta}}=[g(\mathbf{c}_{0}),\nabla g(\mathbf{c}_{0})^{\text{T}}]^{\text{T}}$. The expectation of $\hat{\bm{\theta}}-\bar{\bm{\theta}}$ is given by
$\mathbb{E}\{\hat{\bm{\theta}}-\bar{\bm{\theta}}\}$ $=\mathbb{E}\left\{\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\mathbf{y}\right\}-\bar{\bm{\theta}}$
$=\mathbb{E}\left\{\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\bar{\bm{\theta}}+\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}(\mathbf{e}+\boldsymbol{\xi})\right\}$
$-\bar{\bm{\theta}}$
$=\left(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}}\right)^{-1}\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\mathbf{e}$
$=\frac{1}{M}\left[\begin{array}[]{c}\sum_{m=1}^{M}e(\mathbf{x}_{m},\mathbf{c}_{0})\\ \frac{3}{r_{1}^{2}}\sum_{m=1}^{M}(\mathbf{x}_{m}-\mathbf{c}_{0})\cdot e(\mathbf{x}_{m},\mathbf{c}_{0})\end{array}\right].$
Since $\hat{g}(\mathbf{x})=\hat{\alpha}+\hat{\boldsymbol{\beta}}^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})$ and $g(\mathbf{x})=g(\mathbf{c}_{0})+\nabla g(\mathbf{c}_{0})^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})+e(\mathbf{x},\mathbf{c}_{0})$, we have
$\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}$ $=\left[\begin{array}[]{cc}1&(\mathbf{x}-\mathbf{c}_{0})^{\text{T}}\end{array}\right]\cdot\mathbb{E}\{\hat{\bm{\theta}}-\bar{\bm{\theta}}\}+e(\mathbf{x},\mathbf{c}_{0})$
$=\frac{1}{M}\sum_{m=1}^{M}e(\mathbf{x}_{m},\mathbf{c}_{0})+e(\mathbf{x},\mathbf{c}_{0})$
$+\frac{3}{Mr_{1}^{2}}\sum_{m=1}^{M}(\mathbf{x}-\mathbf{c}_{0})^{\text{T}}(\mathbf{x}_{m}-\mathbf{c}_{0})\cdot e(\mathbf{x}_{m},\mathbf{c}_{0})$
In addition, as $d(\mathbf{x}_{m},\mathbf{c}_{0})\leq r_{1}$ and $d(\mathbf{x},\mathbf{c}_{0})\leq r_{0}$, the Lipschitz condition implies that $|e(\mathbf{x}_{m},\mathbf{c}_{0})|\leq\frac{L_{g}^{2}}{2}d(\mathbf{x}_{m},\mathbf{c}_{0})^{2}\leq L_{g}r_{1}^{2}/2$ and $|e(\mathbf{x},\mathbf{c}_{0})|\leq L_{g}r_{0}^{2}/2$. The bias can be upper bounded as
$\big{|}\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\big{|}$ $\leq\frac{L_{g}r_{1}^{2}}{2}+\frac{L_{g}r_{0}^{2}}{2}+\frac{3}{Mr_{1}^{2}}\cdot Mr_{0}r_{1}\cdot\frac{L_{g}r_{1}^{2}}{2}$
$=\frac{L_{g}}{2}\left(r_{1}^{2}+r_{0}^{2}+3r_{0}r_{1}\right).$ (50)
Finally, substituting (49) and (50) into (48), we have
$\mathbb{E}\left\{\left(\hat{g}(\mathbf{x})-g(\mathbf{x})\right)^{2}\right\}$ $=\mathbb{V}\left\{\hat{g}(\mathbf{x})\right\}+\left(\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\right)^{2}$
$\leq\frac{\sigma^{2}}{M}\left(1+\frac{3r_{0}^{2}}{r_{1}^{2}}\right)$
$+\frac{L_{g}^{2}}{4}\left(r_{1}^{2}+r_{0}^{2}+3r_{0}r_{1}\right)^{2}.$ (51)
Furthermore, if $g(\mathbf{x})\approx g(\mathbf{c}_{0})+\nabla g(\mathbf{x})^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})+L^{\prime}_{g}\|\mathbf{x}-\mathbf{c}_{0}\|^{2}/2$ and $\|\mathbf{x}_{m}-\mathbf{c}_{0}\|^{2}=r_{1}^{2}$, then we have $e(\mathbf{x},\mathbf{c}_{0})=L^{\prime}_{g}\|\mathbf{x}-\mathbf{c}_{0}\|^{2}/2=L_{g}^{\prime}r_{0}^{2}/2$ and $e(\mathbf{x}_{m},\mathbf{c}_{0})=L_{g}^{\prime}r_{1}^{2}/2$ for all $m$. The expression (50) thus simplifies to $\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\approx L_{g}^{\prime}\left(r_{1}^{2}+r_{0}^{2}\right)/2$. Together with the result (49), it leads to the MSE approximation (20).
Appendix E Supplementary Materials 1: Complexity Analysis
As this paper studies trajectory design for UAV online search, the following two metrics are essential for understanding the complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course.
Proposition 3 derives the upper bound of the search trajectory under certain conditions in a balancing problem.
Proposition 3 (Upper bound of trajectory length).
Suppose that the channel gain satisfies $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$. For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the trajectory length $L$ is found to be linear in some geometry parameters of the environment:
$L\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}$ (52)
where $H_{0}$ is the initial search altitude, $R$ is the radius of the equipotential surface given in (11), and $r_{1}$ is the measurement radius.
Proof.
The superposed trajectory proposed in this paper combines the search trajectory ${\bf x}_{\text{s}}(t)$ on the equipotential surface as shown in Fig. 3 with the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$ as shown in Fig. 2.
We first derive the trajectory length of the search trajectory ${\bf x}_{\text{s}}(t)$ on the equipotential surface. According to Proposition 2, the equipotential surface is a sphere centered at $\mathbf{o}=(P_{0}\sum_{k\in\mathcal{K}}\mathbf{u}_{k}-P_{\text{T}}\mathbf{u}_{0})/(KP_{0}-P_{\text{T}})$ with radius $R$ in (11). According to Section IV-B, the search trajectory ${\bf x}_{\text{s}}(t)$ is categorized into the following two phases:
Phase 1: trajectory ${\bf x}_{\text{s}}(t)$ in the full-LOS region. The trajectory in the full-LOS region is an arc with a radius $R$ and a central angle less than $\pi$. Hence, the trajectory ${\bf x}_{\text{s}}(t)$ in the full-LOS region is upper bounded by $\pi R$.
Phase 2: trajectory in the non-full-LOS region. The trajectory in the non-full-LOS region satisfies $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=C$. Therefore, ${\bf x}_{\text{s}}(t)$ in the non-full-LOS region is an arc with a radius smaller than or equal to $H_{0}$ and a central angle less than or equal to $\pi$. Then, the trajectory $\mathbf{x}_{\text{s}}(t)$ in the non-full-LOS region is upper bound by $\pi H_{0}$.
By combining the trajectory length in phase 1 and phase 2, the length of search trajectory $\mathbf{x}_{\text{s}}(t)$ is upper bounded by $\pi(H_{0}+R)$.
Next, we derive the trajectory length of the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$. The spiral measurement trajectory shown in Fig. 2 (b) can be described by the following 3D helix equations:
$\begin{cases}x_{\text{r}1}(t)=\sqrt{2/3}r_{1}\cos(\omega t)\\ x_{\text{r}2}(t)=vt\\ x_{\text{r}3}(t)=\sqrt{2/3}r_{1}\sin(\omega t)\end{cases}$ (53)
where recall that $\omega$ is the angular velocity, $v$ is the velocity in the spiral axis direction, and $r_{1}$ is the measurement radius.
Therefore, the spiral trajectory length denoted as $L_{\text{r}}$ during $t\in[0,T]$ is given by
$L_{\text{r}}$ $=\int_{0}^{T}\sqrt{(-\sqrt{\frac{2}{3}}r_{1}\omega\sin(\omega t))^{2}+(\sqrt{\frac{2}{3}}r_{1}\omega\cos(\omega t))^{2}+v^{2}}\text{d}t$
$=\sqrt{\frac{2r_{1}^{2}\omega^{2}}{3}+v^{2}}T.$ (54)
Finally, we combine the LOS discovery trajectory ${\bf x}_{\text{s}}(t)$ and the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$. Given the upper bound of ${\bf x}_{\text{s}}(t)$ as $\pi(H_{0}+R)$, the total search time is upper bound by $T\leq\pi(H_{0}+R)/v$. According to the design of the spiral trajectory in Section III-E, let $\omega=4\pi/M$ and $v=2/\sqrt{M^{2}-1}$. Then, the total length $L$ of the superposed trajectory is upper bound by
$L$ $\leq\frac{\pi(H_{0}+R)}{v}\sqrt{\frac{2r_{1}^{2}\omega^{2}}{3}+v^{2}}$
$=\pi(H_{0}+R)\sqrt{\frac{8(M^{2}-1)\pi^{2}r_{1}^{2}}{3M^{2}}+1}$
$\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}.$ (55)
Proposition 3 implies that reducing the initial search altitude is likely to reduce the search trajectory length $L$ as it results in a smaller upper bound of $L$. In addition, the radius $R$ of the equipotential surface given in (11) consists of the distance between the user cluster and the BS. Consider a special case where ${\bf u}_{0}=[0,0,0]^{\text{T}}$, $K=1$, $P_{0}\neq P_{\text{T}}$, and $P_{0}>P_{\text{T}}>0$. Then, $R^{2}$ in (11) can be simplified as
$R^{2}=\frac{P_{0}}{P_{0}-P_{\text{T}}}\left(\frac{P_{0}}{P_{0}-P_{\text{T}}}-1\right)d^{2}({\bf u}_{0},{\bf u}_{1}).$ (56)
The equation (56) indicates that the radius $R$ is monotonically increasing w.r.t. the distance $d({\bf u}_{0},{\bf u}_{1})$ between the user and the BS. Thus, the trajectory length is also related to $d({\bf u}_{0},{\bf u}_{1})$. Furthermore, the power allocation also affects the shape of the equipotential surface. For example, when the UAV transmit power $P_{\text{T}}$ is much smaller than the BS transmit power $P_{0}$, the radius $R$ of the equipotential surface in (56) will be quite small.
As the proposed scheme is an online strategy implemented on UAV s, the per-step computational complexity is also derived. Specifically, the computational complexity of the proposed algorithm mainly comes from the local channel construction in Step 3), and the update of UAV positions in Step 4) in Algorithm 1.
  • Computational complexity of Step 3): The local channel construction in Step 3) builds a local channel model for each user using the closed-form formula in equation (17), and thus, the computational complexity is given by $O(KM)$, where recall that $K$ is the number of users and $M$ is the number of measurements used for local channel model construction.
  • Computational complexity of Step 4): The update of UAV positions in Step 4) has a computational complexity of $O((K+N)^{3})$, where $N$ is the number of constraints for resource allocation. Step 4) depends on closed-form formulations in (14), (37), (38), (39), (40), (42), (43) and (44), and the major complexity comes from the matrix inverse operation in (37).
Appendix F Supplementary Materials 2: Proof of Lemma 1
We first state the following two lemmas.
Lemma 2 (Eigenvalue inequality [32, Corollary 4.3.12]).
For a positive definite symmetric matrix $\mathbf{B}\in M_{N}$ and $c\geq 0$, we have
$0<\lambda_{i}(\mathbf{B})\leq\lambda_{i}(\mathbf{B}+c),\quad i=1,2,\dots,N$
with the equality for some $i$ if and only if $c=0$.
Lemma 3 (Trace inequality [32, Equation 2.4.2.1]).
For a positive definite symmetric matrix $\mathbf{B}\in M_{N}$ and $c\geq 0$, we have
$\text{tr}\left\{\left(\mathbf{B}+c\right)^{-1}\right\}=\sum_{i=1}^{N}\lambda_{i}^{-1}(\mathbf{B}+c)\leq\sum_{i=1}^{N}\lambda_{i}^{-1}(\mathbf{B})=\text{tr}\left\{\mathbf{B}^{-1}\right\}$
with the equality if and only if $c=0$.
Denote $\mathbf{A}_{N-i}\in M_{N-i}$ as a submatrix of $\mathbf{A}$ by removing the first $i$ rows and the first $i$ columns. Denote $\mathbf{a}_{-i}$ as a column vector that contains the lower off-diagonal elements of the $i$th column of $\mathbf{A}$. Thus,
$\mathbf{A}_{N}=\left[\begin{array}[]{cc}a_{11}&\mathbf{a}_{-1}\\ \mathbf{a}_{-1}^{\text{T}}&\mathbf{A}_{N-1}\end{array}\right],\,\mathbf{A}_{N-1}=\left[\begin{array}[]{cc}a_{22}&\mathbf{a}_{-2}\\ \mathbf{a}_{-2}^{\text{T}}&\mathbf{A}_{N-2}\end{array}\right].$
Using the block matrix inverse, we have
$\mathbf{A}_{N}^{-1}$
$=\left[\begin{array}[]{cc}(a_{11}-\mathbf{a}_{-1}^{\text{T}}\mathbf{A}_{N-1}\mathbf{a}_{-1})^{-1}&*\\ *&(\mathbf{A}_{N-1}-\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1})^{-1}\end{array}\right].$
Since $\mathbf{A}$ is positive definite, so are $a_{11}$ and $\mathbf{A}_{N-1}$. Thus, $\mathbf{a}_{-1}^{\text{T}}\mathbf{A}_{N-1}\mathbf{a}_{-i}\geq 0$ and $\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1}\geq 0$ for all $\mathbf{a}_{-1}$ with equalities both at $\mathbf{a}_{-1}=\mathbf{0}$.
Thus, let $\mathbf{B}=\mathbf{A}_{N-1}-\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1}$, and $c=\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1}$ where $\mathbf{B}$ must be positive definite because $\mathbf{A}_{N}$ is positive definite, so as $\mathbf{A}_{N}^{-1}$ and $\mathbf{A}_{N-1}-\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1}$. Additionally, since $a_{11}>0$, $c\geq 0$ with the equality if and only if $\mathbf{a}_{-1}=\mathbf{0}$. Given Lemma 2 and Lemma 3, we have $\text{tr}\left\{\mathbf{B}^{-1}\right\}\geq\text{tr}\{\left(\mathbf{B}+c\right)^{-1}\}$ with equality when $\mathbf{a}_{-1}=\mathbf{0}$. As a result,
$\text{tr}\{\mathbf{A}_{N}^{-1}\}$ $\geq a_{11}^{-1}+\text{tr}\{\left(\mathbf{A}_{N-1}-\mathbf{a}_{-1}^{\text{T}}a_{11}\mathbf{a}_{-1}\right)^{-1}\}$
$\geq a_{11}^{-1}+\text{tr}\{\mathbf{A}_{N-1}^{-1}\}$
with the equality when $\mathbf{a}_{-1}=\mathbf{0}$.
Similarly, we have
$\text{tr}\{\mathbf{A}_{N-1}^{-1}\}$ $\geq a_{22}^{-1}+\text{tr}\{\mathbf{A}_{N-2}^{-1}\},\dots$
$\vdots$
$\text{tr}\{\mathbf{A}_{2}^{-1}\}$ $\geq a_{N-1,N-1}^{-1}+\text{tr}\{a_{N,N}^{-1}\}.$
Combining the above inequalities, one can obtain that $\text{tr}\{\mathbf{A}^{-1}\}\geq\sum_{i=1}^{N}a_{ii}^{-1}$ with the equality achieved for $a_{ij}=0$, $\forall i\neq j$.

### v4

@@
 An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel.
-Numerical results on real 3D city maps demonstrate that the proposed scheme achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search.
+Numerical results on real 3D city maps demonstrate that the proposed scheme achieves over $94\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search.
 Index Terms:
-UAV trajectory design, unknown user locations, unknown channels, unknown environment, perturbation theory
+UAV placement, unknown user locations, perturbation theory, trajectory design, channel map construction
 I Introduction
-In dense urban environments, short-wavelength signals such as millimeter wave (mmW), Terahertz (THz), and optical waves experience significant attenuation due to obstructions like buildings and other structures [1].
-This attenuation is particularly problematic for sensing performance, which depends heavily on maintaining line-of-sight (LOS) links between targets and transceivers [2, 3].
-Deploying a low-altitude unmanned aerial vehicle (UAV) presents a promising solution for establishing LOS links for users in deep shadowed areas, as UAV s operate at relatively high altitudes with the flexibility of 3D mobility.
-However, optimizing UAV deployment remains challenging due to the arbitrarily complex terrain topology, which complicates modeling and predicting LOS conditions.
-Previous works have circumvented this challenge by assuming distance-dependent pure LOS channel models [4, 5], or probabilistic channel models [6, 7, 8, 9].
-While these simplified models allow for coarse analysis, they often compromise performance by neglecting the actual LOS conditions.
-Some state-of-the-art works have developed blockage-aware channel models [10, 11] to enhance UAV placement strategies.
-For instance, the authors in [10] approximated buildings as polyhedrons and derived explicit expressions to determine the LOS status of a 3D position using an offline city map.
-However, constructing geometric maps or radio maps incurs substantial costs, and polyhedral approximation, while useful, remains a coarse and computationally intensive method, making it challenging to scale to larger areas.
-In addition, most existing studies [4, 5, 6, 7, 8, 9, 10, 11, 12] assume perfect knowledge of user locations and channel models.
-However, in practical scenarios, accurate user locations or precise channel models are typically unavailable due to the location privacy of the users and the uncertain non-line-of-sight (NLOS) condition.
-Some recent research efforts have considered first estimating user locations and channel parameters before optimizing UAV service positions [13, 14, 15].
-Nevertheless, localization under potential NLOS conditions is already a challenging problem, and the performance of UAV placement is highly dependent on the accuracy of user localization.
-Assuming the availability of 3D maps, the work in [15] employed the Fisher information metric to design an online trajectory for joint path-loss parameter estimation and user localization, although the global optimality and complexity of the approach remain unclear.
-With known user locations and maps, the authors in [13] utilized dynamic programming to optimize UAV trajectory for channel parameter learning, employing map compression to estimate LOS probability and assist in UAV placement.
-However, the effectiveness of the placement heavily depends on channel learning and LOS probability estimation.
-In the absence of prior knowledge of channels and user locations, a model-free approach presented in [14] relied on signal-to-interference-and-noise ratio (SINR) measurements and a 3D map to account for blockages and scatterings, applying Q-learning to optimize UAV placement.
-Despite this, the action space was limited to four orthogonal horizontal directions, significantly reducing the likelihood of identifying the optimal UAV placement in a 3D environment.
-This paper investigates the UAV trajectory design to provide communication and sensing services for a cluster of ground users at unknown locations, while also maintaining backhaul connectivity with a remote base station (BS).
-The goal is to develop an optimization framework for an efficient online search of the best UAV position for a broad class of communication problems and sensing problems.
-The challenges arise from the need to operate without precise information: user locations are unknown due to privacy concerns and potential NLOS conditions, the channel model parameters are difficult to determine due to environmental sensitivity, and the lack of accurate 3D maps further complicates the UAV placement in dense urban environments.
-In this work, we develop three key techniques to address the aforementioned challenges.
-First, we introduce a search method on an equipotential surface using trajectories derived from perturbation theory.
-The equipotential surface represents a region where the sensing and communication performance for users is balanced with the capacity of the UAV-BS backhaul link.
-The UAV’s proximity to the equipotential surface can be assessed by measuring the signal-to-noise ratio (SNR), without requiring knowledge of user locations or channel parameters.
-Although the exact expression of the equipotential surface remains elusive, perturbation theory guides the UAV search direction while maintaining its position on this surface.
-Second, we propose a local construction of channel models for each user within the LOS regime using local polynomial regression.
-This approach allows for a relatively simple construction of local LOS models without requiring user location or propagation distance.
-Third, we develop a search strategy on the equipotential surface that leverages two universal properties: upward invariance and colinear invariance of LOS regions over almost any terrain structure.
-Prior studies have shown that, in certain cases, this strategy can achieve an $\epsilon$-optimal solution globally in 3D space with a trajectory length proportional to the search radius [16, 17].
-However, previous work [17] assumed knowledge of user locations and channel model parameters.
-Contributions of this paper are as follows:
-  • We study UAV trajectory design for serving a cluster of users at unknown locations with unknown channels.
-Specifically, we develop our search strategy based on perturbation theory and two universal properties, i.e., upward invariance and colinear invariance, of the LOS regions over almost all terrain structure.
-  • We provide an optimal design for a class of UAV search trajectories to balance the estimation for user channels and the exploration for better channel quality.
-This also includes deriving the upper bound of the mean squared error (MSE) for the locally constructed channel gain based on an optimal measurement pattern and determining the optimal radius of the spiral trajectory for minimizing the MSE of channel gain estimation.
-  • We demonstrate through numerical results that the normalized channel gain construction error is on the order of $10^{-2}$ without knowing user locations or propagation distances.
-Furthermore, with a $3$ kilometer search, the proposed scheme can achieve over $95\%$ performance of that from a 3D exhaustive search over a dense urban area.
+In dense urban environments, short-wavelength signals like millimeter wave (mmW), Terahertz (THz), and optical waves suffer from severe attenuation due to signal blockage from buildings and other structures [1].
+In particular, sensing performance heavily relies on line-of-sight (LOS) links between targets and transceivers, which are usually disrupted in urban environments [2, 3].
+Deploying a low-altitude unmanned aerial vehicle (UAV) is a promising solution for providing LOS links for users in deep shadow since the UAV operates at a relatively high altitude with 3D mobility.
+Nevertheless, the optimization of UAV deployment poses challenges due to the arbitrarily complex terrain topology, which complicates modeling and predicting LOS conditions.
+Previous works bypassed this challenge by assuming distance-dependent pure LOS channel models [4, 5], or probabilistic channel models [6, 7, 8, 9].
+These simplified channel models enable a coarse analysis but sacrifice the performance by ignoring the actual LOS condition.
+Some state-of-the-art works formulated blockage-aware channel models [10, 11] to assist UAV placement.
+For example, the authors in [10] approximated buildings as polyhedrons, and derived explicit expressions to determine the LOS status of a 3D position using an offline city map.
+However, the performance of these methods highly relies on the accuracy of the model of the environment.
+In addition, most existing works [4, 5, 6, 7, 8, 9, 10, 11, 12] assume perfect knowledge of user locations and channel models.
+Nevertheless, in practical scenarios, accurate user locations and precise channel models are typically unavailable due to the location privacy of users and the uncertain non-line-of-sight (NLOS) conditions.
+Some recent works considered first estimating user locations and channel parameters before optimizing UAV service positions [13, 14, 15].
+However, localization itself under possibly NLOS conditions is already a challenging problem and the performance of UAV placement highly depends on the localization accuracy of the users.
+Assuming the availability of 3D maps, the work [15] employed the fisher information metric to design an online trajectory for joint path-loss parameter estimation and user localization, but the global optimality and complexity remain unclear.
+With known user locations and maps, the work [13] utilized dynamic programming to optimize UAV trajectory for channel parameters learning and employed map compression to estimate the LOS probability for UAV positions.
+However, the placement performance relies on channel learning and LOS probability estimation.
+Without the prior knowledge of channels and user locations, a model-free approach in [14] relied on signal-to-interference-and-noise ratio (SINR) measurements and a 3D map to account for blockage and scatters, and the authors applied Q-learning to optimize UAV placement.
+However, the action space is restricted to four orthogonal horizontal directions on a 2D plane, and there is little theoretical guarantee on the performance.
+This paper investigates the UAV trajectory design to provide communication and sensing services for a cluster of ground users at unknown locations, while the UAV maintains backhaul connectivity with a remote base station (BS).
+The goal is to provide an optimization framework for an efficient online search of the best UAV position for a broad class of communication problems and sensing problems.
+The challenges arise from the need to operate without precise information, where the user locations and the channel models are not available at the UAV, and the UAV has to make search decisions and adjust the search trajectory on-the-fly based on the online measurements.
+In this work, we develop three techniques to resolve the above challenges.
+First, we propose to search on an equipotential surface using trajectories developed from perturbation theory.
+The equipotential surface is the region of UAV positions where the sensing and communication performance for the users is balanced with capacity of the UAV-BS backhaul link.
+The status whether the UAV is on or off the equipotential surface can be quantified by measuring the signal-to-noise ratio (SNR), without requiring knowledge of user locations or channel parameters.
+While an explicit expression of the equipotential surface is not available, we employ perturbation theory to develop a search strategy so that the search trajectory remains on the equipotential surface.
+Second, we propose to locally construct a channel map for each user within the LOS regime using local polynomial regression.
+This approach allows for a relatively simple construction of nonparametric channel models that are capable of capturing actual signal attenuations due to blockage from the possibly unknown propagation environment.
+Third, we develop our search strategy on the equipotential surface exploiting two universal properties: upward invariance and colinear invariance of LOS regions over almost all terrain structure.
+Prior studies have shown that, in certain cases, this strategy can achieve an $\epsilon$-optimal solution globally in 3D space with a trajectory length linear of the search radius [16, 17].
+Yet, the prior work [16, 17] assumed knowledge of user locations and channel model parameters.
+The contributions of this paper are summarized as follows:
+  • We analytically show that the equipotential surface is a sphere for a class of sensing and communication problems, where the sphere parameters depend on the user distribution and the power budget.
+  • We develop a class of spiral trajectories to simultaneously construct the local LOS channels and search on the equipotential surface.
+An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel.
+  • We demonstrate that the normalized channel gain construction error is on the order of $10^{-2}$ without knowing the user location or the propagation distance.
+With a $3$-kilometer search, the proposed scheme can achieve over $94\%$ performance of that from a 3D exhaustive search for a UAV-assisted multiuser sensing and communication problem over a dense urban area.
 The remaining part of the paper is organized as follows: Section II introduces the system model and formulates a UAV-assisted sensing and communication problem.
-Section III discusses the geometric properties of the equipotential surface and local channel model construction.
+Section III discusses the geometric properties of the equipotential surface and local channel map construction.
 Section IV outlines the LOS discovery trajectory and proposes a superposed trajectory for optimal LOS position search.
@@
 The sets of users are denoted as $\mathcal{K}\triangleq\{1,2,\dots,K\}$.
-Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{3}:x_{3}\geq H_{\text{min}}\}$ be the feasible region of the UAV positions.
+Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{3}:x_{3}\geq H_{\text{min}}\}$ be the feasible region of UAV positions where $H_{\text{min}}$ is the minimum flight height.
 While signals of ground users are likely blocked by buildings, we denote $\mathcal{D}_{k}\text{$\subseteq$}\mathcal{X}$ as the region of UAV positions such that there is an LOS link between the UAV at position $\mathbf{x}\in\mathcal{D}_{k}$ and user $k$ at an unknown position $\mathbf{u}_{k}$.
@@
 $\quad H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0,n=1,2,\dots,N$
-where $f_{0}(g_{0}(\mathbf{x}))$ represents the objective of the BS-UAV link under the LOS condition, $F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ represents the objective of the UAV-user links under LOS, and $H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0$ are the corresponding constraints for the resource allocation.
-Problem 5 is non-convex due to the LOS constraint on UAV positions and the fact that the blockage can have an arbitrary shape.
+where $f_{0}(g_{0}(\mathbf{x}))$ represents the objective of the BS-UAV link under the LOS condition, $F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ represents the objective of the UAV-user links under LOS, and $H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0$ for $n=1,2,\dots,N$ are the corresponding constraints for the resource allocation.
+Problem $\mathscr{P}$ is non-convex due to the LOS constraint on UAV positions and the fact that the blockage can have an arbitrary shape.
 Such a general formulation captures many typical applications for communication and sensing, with two examples illustrated as follows.
 II-C1 Balancing problem
-Consider to deploy a UAV to offer sensing and communication services for a cluster of sensing targets and communication users while establishing a LOS relay link with a remote BS.
+Consider to deploy a UAV to offer sensing and/or communication services for a cluster of sensing targets and/or communication users while establishing a LOS relay link with a remote BS.
 Specify $p_{k}$ as the power allocation from the UAV to user $k$.
@@
 For sake of brevity, let $N_{0}=1$ in this paper.
-For a sensing task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as weighted SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{s}}p_{k}g_{k}(\mathbf{x})$ where $\mu_{\text{s}}$ is a weight of the sensing task [23, 24].
+For a sensing task involving estimation, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as weighted SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{s}}p_{k}g_{k}(\mathbf{x})$ where $\mu_{\text{s}}$ is a weight of the sensing task [23, 24].
 For a communication task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as channel capacity, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{c}}\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$ where $\mu_{\text{c}}$ is a weight of the communication task.
@@
 $\quad\sum_{k\in\mathcal{K}}p_{k}\leq P_{\text{T}}.$
-While the balancing problem is non-convex and the sum-rate problem is convex in the resource allocation variable $\mathbf{p}$, the whole problem $\mathscr{P}$ that jointly optimizes the UAV location $\mathbf{x}$ is non-convex.
+While the example formulation (6) is non-convex in the power allocation variable $\mathbf{p}$ and the problem (7) is convex in $\mathbf{p}$, both cases can be handled in a same way in our proposed algorithm framework.
 In addition, as the channels and the LOS regions $\mathcal{D}_{k}$ are not available before exploring near location $\mathbf{x}$, the UAV needs to design an online trajectory to explore the LOS opportunity, measure the channel quality, and optimize for the system performance.
@@
 In the rest of this section, we develop methods and trajectory $\mathbf{x}_{\text{r}}(t)$ to locally construct $\bm{g}(\mathbf{x})$ and its gradient $\mathbf{G}(\mathbf{x})$ at the neighborhood of $\mathbf{x}_{\text{s}}(t)$.
-III-C Construction of a Local Channel Model
+III-C Construction of a Local Channel Map
 It is well-known that directly estimating a global propagation model $g_{k}(\mathbf{x})$ is difficult, as it leads to nonlinear regression.
-Instead, for each node $k$, we adopt a linear model $\hat{g}(\mathbf{x})$ to locally approximate the LOS channel gain $g_{k}(\mathbf{x})$ for the LOS region $\mathbf{x}\in\mathcal{D}_{k}$ in the neighborhood of $\mathbf{x}=\mathbf{c}_{0}$:
+Instead, for each node $k$, we adopt a linear model $\hat{g}(\mathbf{x})$ to locally approximate the LOS channel map $g_{k}(\mathbf{x})$ for the LOS region $\mathcal{D}_{k}$ in the neighborhood of $\mathbf{x}=\mathbf{c}_{0}$:
 $\hat{g}(\mathbf{x})=\alpha+\boldsymbol{\beta}^{\text{T}}(\mathbf{x}-\mathbf{c}_{0})$ (16)
@@
 One can easily verify that if we sample at $t=m-1/2$, i.e., $\mathbf{x}_{m}=\mathbf{x}(m-1/2)$, for $m=1,2,\dots,M$, then conditions (i)–(iii) in Theorem 1 are satisfied.
-As a result, the MSE of the locally reconstructed channel $\hat{g}_{k}(\mathbf{x})$ is upper bound by (19).
+As a result, the MSE of the locally reconstructed channel $\hat{g}_{k}(\mathbf{x})$ is upper bounded by (19).
 Figure 2: (a) An alternating spiral trajectory (orange dots) that satisfies conditions (i)–(iii) in Theorem 1.
@@
 To circumvent this issue, a transition phase $t\in(t_{1},t_{1}+\tau)$ is needed, where without altering $\mathbf{x}_{\text{s}}(t)$, i.e., $\dot{\mathbf{x}}_{\text{s}}=0$, we smoothly switch $\dot{\mathbf{x}}_{\text{s}}$ from $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})$ to $\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})$ using a linear transition
-$\dot{\mathbf{x}}_{\text{t}}=\frac{\tau-t+t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})+\frac{t-t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+}),\quad t\in(t_{1},t_{1}+\tau)$ (43)
+$\dot{\mathbf{x}}_{\text{t}}=\frac{\tau-t+t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})+\frac{t-t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+}),\>t\in(t_{1},t_{1}+\tau)$ (43)
 which yields $\ddot{\mathbf{x}}_{\text{t}}=\frac{1}{\tau}(\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+})-\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-}))$.
@@
 A sample implementation is summarized in Algorithm 1.
-As Algorithm 1 is an online search scheme, the following two metrics are essential for understanding its complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course.
-Specifically, given a common realization of channel models and a specified objective function for the balancing problem in Proposition 2, the trajectory length $L$ is found to be linear in some geometry parameters of the environment, i.e.,
+IV-C3 Complexity Analysis
+As Algorithm 1 is an online search scheme, we investigate two different metrics for an understanding of its complexity: the trajectory length of the UAV and the per-step computational complexity every time the UAV adjusts its course.
+The following proposition shows the trajectory length is linear in the radius of the equipotential surface under certain conditions.
+Proposition 3 (Upper bound of trajectory length).
+Suppose that $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$.
+For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the trajectory length $L$ is upper bounded as
 $L\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}$ (45)
-where $H_{0}$ is the initial search altitude, $R$ is the radius of $\mathcal{S}$ given in (11), and $r_{1}$ is the measurement range defined in Section III.
-Recall that $K$ is the number of users, $M$ is the number of measurements used for local channel model construction, and $N$ is the number of constraints for resource allocation.
-The per-step computational complexity is found as $O(KM+(K+N)^{3})$, which is numerically found to be handled in milliseconds.
-For detailed complexity derivations, see Appendix E: Supplementary Material 1.
+where $H_{0}$ is the initial search altitude, $R$ is the radius of the equipotential surface given in (11), and $r_{1}$ is the measurement range.
+Proof.
+See Appendix E in [27]. ∎
+Proposition 3 suggests that reducing the initial search altitude decreases the upper bound of the trajectory length $L$.
+Additionally, the trajectory length $L$ is related to the radius $R$ of the equipotential surface.
+In a special case where ${\bf u}_{0}=[0,0,0]^{\text{T}}$, $K=1$, and $P_{0}>P_{\text{T}}>0$, $R^{2}$ in (11) simplifies to $R^{2}=P_{0}/(P_{0}-P_{\text{T}})\left(P_{0}/(P_{0}-P_{\text{T}})-1\right)d^{2}({\bf u}_{0},{\bf u}_{1})$, which indicates that $R$ is linear in $d({\bf u}_{0},{\bf u}_{1})$.
+Thus, the worst-case search trajectory length is linear in the BS-user distance $d({\bf u}_{0},{\bf u}_{1})$.
+Recall that $K$ is the number of users, $M$ is the number of measurements used for local channel construction, and $N$ is the number of constraints for resource allocation in problem (22).
+The computational complexity of Steps 3 and 4 in Algorithm 1 is found as $O(KM+(K+N)^{3})$.
+See Appendix F in [27] for a detailed derivation of the computational complexity.
 Find a full-LOS initial position $\mathbf{x}_{0}\in\mathcal{S}\cap\tilde{\mathcal{D}}$. Denote $\mathbf{x}_{\text{r}}(t)=r[\cos(\omega t),\,0,\,\sin(\omega t)]^{\text{T}}$.
@@
     Local channel construction: Collect channel measurements for each time slot $\triangle t$.
-Construct the local channel $\hat{g}_{k}(\mathbf{x})$ and obtain the estimate for $\nabla g_{k}(\mathbf{x})$ for each user based on model (16) with parameter (17) estimated from the past $M$ LOS measurements.
+Construct the local channel $\hat{g}_{k}(\mathbf{x})$ and obtain the estimation of $\nabla g_{k}(\mathbf{x})$ for each user based on model (16) with parameters (17) estimated from the past $M$ LOS measurements.
   4. 4.
@@
 It is assumed that the mmW beam alignment has been done for every UAV position.
-For a balancing application, one UAV is placed to provide detection services for eight sensing targets and maintains a backhaul communication link with a BS simultaneously.
-Particularly, the objective functions for sensing are specified as SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=p_{k}g_{k}(\mathbf{x})/(\bar{N}_{0}W)$ where the noise power density $\bar{N}_{0}$ is set as $-150$ dBm/Hz, and the bandwidth $W$ is set as $1$ GHz.
+For a balancing application, one UAV is placed to provide location estimation services for $8$ sensing targets and maintain a backhaul communication link with a BS simultaneously.
+Particularly, the objective functions for sensing are specified as SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=p_{k}g_{k}(\mathbf{x})/(\bar{N}_{0}W)$ where the noise power spectral density $\bar{N}_{0}$ is set as $-150$ dBm/Hz, and the bandwidth $W$ is set as $1$ GHz.
 Figure 4: Map A (left) is a sparse commercial area, and map B (right) is a dense residential area in Beijing, China.
@@
 Map B 1.70 1.84 2.08 2.17 2.24 2.30
-Figure 7: System capacity versus transmit power of UAV
+Figure 7: System capacity versus transmit power of UAV in a sum-rate application
 Table II summarizes the system capacity, i.e., $\min\{f_{0}(g_{0}(\mathbf{x})),\sum_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k}\text{)}\}\}$, of different schemes on two maps.
@@
 Thus, gathering a sufficient number of LOS measurements in a local area for channel estimation becomes more challenging, resulting in increased construction error of the equipotential surface.
-Figure 8: System capacity versus number of users
-Figure 9: CDF versus $\text{SNR}=\min_{k\in\mathcal{K}}\{p_{k}^{*}g_{k}(\mathbf{x})/N_{0}\}$ in a UAV-assisted sensing and communication application.
+Figure 8: System capacity versus number of users in a sum-rate application
+Figure 9: CDF versus SNR of the weakest sensing link in a balancing application.
+In Fig.
+9, we examine the proposed scheme in a balancing application.
+As observed, the proposed scheme substantially increases the SNR of the weakest sensing link over the statistical geometry scheme.
+In particular, it is observed that the cumulative distribution function (CDF) curve of the proposed scheme closely coincides with that of the exhaustive 3D search on map B, numerically confirming that the proposed scheme can find the near-optimal solution in complex urban environments.
+Fig.
+10 shows two examples of convergence process of the proposed scheme.
+The velocity of the UAV is set as $10$ meters per second.
+It is observed that the system capacity increases with the search time because the proposed scheme searches in the direction of increasing the objective value in the full-LOS region or maintaining the current objective value in the non-full-LOS region.
+Define the convergence time as the search time at which the system performance no longer improves, in which case, we consider that the best UAV position is discovered.
+Fig.
+11 summarizes the convergence time versus number of users.
+It is observed that the convergence time decreases with the increase of number of users $K$.
+The reason is that the radius of the equipotential surface decreases with the increase of $K$, which is analytically shown in Proposition 2.
+Since the maximum height of the equipotential surface is smaller, the trajectory length will be shorter as it takes less time to reach the minimum flight height.
+Table III summarizes the average trajectory lengths required by the four online schemes on two maps.
+Notably, the genius-aided scheme with known user locations and channel models requires only a few hundred meters to approximate a near-optimal solution, confirming the efficiency of the LOS discovery strategy proposed in Section IV-B.
+The proposed scheme demands a longer trajectory of approximately $3$ kilometers due to its reliance on spiral trajectories for data collection during the search process.
+Despite this, the proposed scheme still demonstrates considerable efficacy, significantly reducing search complexity compared to the exhaustive search schemes.
+Figure 10: Convergence process: system capacity versus search time.
+Figure 11: The search time to discover the best UAV position versus the number of users. The shaded area represents a possible search time realization during our $2000$ simulations.
 Table III: Comparison of Average Trajectory Length [kilometer]
@@
 Map B 0.2428 3.051 1920 42240
-In Fig.
-9, we examine the proposed scheme in a balancing application.
-As observed, the proposed scheme substantially increases the SNR of the weakest sensing link over the statistical geometry scheme.
-In particular, it is observed that the cumulative distribution function (CDF) curve of the proposed scheme closely coincides with that of the exhaustive 3D search on map B, numerically confirming that the proposed scheme can find the near-optimal solution in complex urban environments.
-Table III summarizes the average trajectory lengths required by the four online schemes on two maps.
-Notably, the genius-aided scheme with known user locations and channel models requires only a few hundred meters to approximate a near-optimal solution, confirming the efficiency of the LOS discovery strategy proposed in Section IV-B.
-The proposed scheme demands a longer trajectory of approximately $3$ kilometers due to its reliance on spiral trajectories for data collection during the search process.
-Despite this, the proposed scheme still demonstrates considerable efficacy, significantly reducing search complexity compared to the exhaustive search schemes.
 VI Conclusion
-In this paper, we developed an efficient online trajectory for UAV placement without prior knowledge of user locations, channel model parameters, or terrain structure.
-We analytically characterized the equipotential surface and proposed an LOS discovery strategy on it, utilizing perturbation theory to guide the UAV search direction.
-Furthermore, we introduced spiral trajectories to construct local channel models in the LOS regime using local polynomial regression, independent of user locations.
+This paper developed an efficient online trajectory for optimal UAV placement without prior knowledge of user locations, channel model parameters, and terrain structure.
+We analytically characterized the equipotential surface and proposed an LOS discovery trajectory on it, utilizing perturbation theory to guide the UAV search direction.
+Additionally, we developed a class of spiral trajectories to construct a local channel map in the LOS regime using local polynomial regression without depending on user locations.
 After deriving the optimal measurement pattern, we minimized the MSE of the locally estimated channel gain and determined the optimal measurement range.
-Experimental results on real urban maps demonstrated that our approach achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search, offering a practical solution for UAV-assisted sensing and communication in complex environments.
+Experimental results on real urban maps demonstrated that our approach achieves over $94\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search in a complex environment.
 Appendix A Proof of Proposition 1
@@
 Proof.
-See Appendix F: Supplementary Material 2. ∎
+See Appendix G in [27]. ∎
 According to [25, 26], the variance of $\boldsymbol{\theta}^{\text{(e)}}$ is given by $\text{tr}\{\sigma^{2}(\tilde{\boldsymbol{\mathbf{X}}}^{\text{T}}\tilde{\boldsymbol{\mathbf{X}}})^{-1}\}$.
@@
 $\mathbb{E}\left\{\left(\hat{g}(\mathbf{x})-g(\mathbf{x})\right)^{2}\right\}$ $=\mathbb{V}\left\{\hat{g}(\mathbf{x})\right\}+\left(\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\right)^{2}$ (48)
-Next, the bounds of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ and $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$ are derived as below.
+Next, the bounds of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ and $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$ are derived as follows.
 1) Derivation of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$: According to the proof of Theorem 1, the variances of $\hat{\alpha}$ and $\hat{\boldsymbol{\beta}}$ are given by $\mathbb{V}\{\hat{\alpha}\}=\sigma^{2}/M$, and $\mathbb{V}\{\hat{\beta}_{j}\}=3\sigma^{2}/Mr_{1}^{2}$.
@@
 In addition, as $d(\mathbf{x}_{m},\mathbf{c}_{0})\leq r_{1}$ and $d(\mathbf{x},\mathbf{c}_{0})\leq r_{0}$, the Lipschitz condition implies that $|e(\mathbf{x}_{m},\mathbf{c}_{0})|\leq\frac{L_{g}^{2}}{2}d(\mathbf{x}_{m},\mathbf{c}_{0})^{2}\leq L_{g}r_{1}^{2}/2$ and $|e(\mathbf{x},\mathbf{c}_{0})|\leq L_{g}r_{0}^{2}/2$.
-The bias can be upper bounded as
+The bias is upper bounded as
 $\big{|}\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}\big{|}$ $\leq\frac{L_{g}r_{1}^{2}}{2}+\frac{L_{g}r_{0}^{2}}{2}+\frac{3}{Mr_{1}^{2}}\cdot Mr_{0}r_{1}\cdot\frac{L_{g}r_{1}^{2}}{2}$
@@
 Together with the result (49), it leads to the MSE approximation (20).
-Appendix E Supplementary Materials 1: Complexity Analysis
-As this paper studies trajectory design for UAV online search, the following two metrics are essential for understanding the complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course.
-Proposition 3 derives the upper bound of the search trajectory under certain conditions in a balancing problem.
-Proposition 3 (Upper bound of trajectory length).
-Suppose that the channel gain satisfies $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$.
-For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the trajectory length $L$ is found to be linear in some geometry parameters of the environment:
-$L\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}$ (52)
-where $H_{0}$ is the initial search altitude, $R$ is the radius of the equipotential surface given in (11), and $r_{1}$ is the measurement radius.
-Proof.
+Supplementary Materials
+Appendix E Proof of Proposition 3
 The superposed trajectory proposed in this paper combines the search trajectory ${\bf x}_{\text{s}}(t)$ on the equipotential surface as shown in Fig.
@@
 Therefore, ${\bf x}_{\text{s}}(t)$ in the non-full-LOS region is an arc with a radius smaller than or equal to $H_{0}$ and a central angle less than or equal to $\pi$.
-Then, the trajectory $\mathbf{x}_{\text{s}}(t)$ in the non-full-LOS region is upper bound by $\pi H_{0}$.
+Then, the trajectory $\mathbf{x}_{\text{s}}(t)$ in the non-full-LOS region is upper bounded by $\pi H_{0}$.
 By combining the trajectory length in phase 1 and phase 2, the length of search trajectory $\mathbf{x}_{\text{s}}(t)$ is upper bounded by $\pi(H_{0}+R)$.
@@
 Finally, we combine the LOS discovery trajectory ${\bf x}_{\text{s}}(t)$ and the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$.
-Given the upper bound of ${\bf x}_{\text{s}}(t)$ as $\pi(H_{0}+R)$, the total search time is upper bound by $T\leq\pi(H_{0}+R)/v$.
+Given the upper bound of ${\bf x}_{\text{s}}(t)$ as $\pi(H_{0}+R)$, the total search time is upper bounded by $T\leq\pi(H_{0}+R)/v$.
 According to the design of the spiral trajectory in Section III-E, let $\omega=4\pi/M$ and $v=2/\sqrt{M^{2}-1}$.
-Then, the total length $L$ of the superposed trajectory is upper bound by
+Then, the total length $L$ of the superposed trajectory is upper bounded by
 $L$ $\leq\frac{\pi(H_{0}+R)}{v}\sqrt{\frac{2r_{1}^{2}\omega^{2}}{3}+v^{2}}$
@@
 $\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}.$ (54)
-Proposition 3 implies that reducing the initial search altitude is likely to reduce the search trajectory length $L$ as it results in a smaller upper bound of $L$.
-In addition, the radius $R$ of the equipotential surface given in (11) consists of the distance between the user cluster and the BS.
-Consider a special case where ${\bf u}_{0}=[0,0,0]^{\text{T}}$, $K=1$, $P_{0}\neq P_{\text{T}}$, and $P_{0}>P_{\text{T}}>0$.
-Then, $R^{2}$ in (11) can be simplified as
-$R^{2}=\frac{P_{0}}{P_{0}-P_{\text{T}}}\left(\frac{P_{0}}{P_{0}-P_{\text{T}}}-1\right)d^{2}({\bf u}_{0},{\bf u}_{1}).$ (56)
-The equation (56) indicates that the radius $R$ is monotonically increasing w.r.t.
-the distance $d({\bf u}_{0},{\bf u}_{1})$ between the user and the BS.
-Thus, the trajectory length is also related to $d({\bf u}_{0},{\bf u}_{1})$.
-Furthermore, the power allocation also affects the shape of the equipotential surface.
-For example, when the UAV transmit power $P_{\text{T}}$ is much smaller than the BS transmit power $P_{0}$, the radius $R$ of the equipotential surface in (56) will be quite small.
-As the proposed scheme is an online strategy implemented on UAV s, the per-step computational complexity is also derived.
-Specifically, the computational complexity of the proposed algorithm mainly comes from the local channel construction in Step 3), and the update of UAV positions in Step 4) in Algorithm 1.
-  • Computational complexity of Step 3): The local channel construction in Step 3) builds a local channel model for each user using the closed-form formula in equation (17), and thus, the computational complexity is given by $O(KM)$, where recall that $K$ is the number of users and $M$ is the number of measurements used for local channel model construction.
-  • Computational complexity of Step 4): The update of UAV positions in Step 4) has a computational complexity of $O((K+N)^{3})$, where $N$ is the number of constraints for resource allocation.
-Step 4) depends on closed-form formulations in (14), (37), (38), (39), (40), (42), (43) and (44), and the major complexity comes from the matrix inverse operation in (37).
-Appendix F Supplementary Materials 2: Proof of Lemma 1
+Appendix F Per-step Computational Complexity Analysis
+The computational complexity of the proposed algorithm mainly comes from the local channel construction in Step 3, and the update of UAV positions in Step 4 in Algorithm 1.
+  • Computational complexity of Step 3: The local channel construction in Step 3 builds a local channel model for each user using the closed-form formula in equation (17), and thus, the computational complexity is given by $O(KM)$, where recall that $K$ is the number of users and $M$ is the number of measurements used for local channel model construction.
+  • Computational complexity of Step 4: The update of UAV positions in Step 4 has a computational complexity of $O((K+N)^{3})$, where $N$ is the number of constraints for resource allocation.
+Step 4 depends on closed-form formulations in (14), (37), (38), (39), (40), (42), (43) and (44), and the major complexity comes from the matrix inverse operation in (37).
+Appendix G Proof of Lemma 1
 We first state the following two lemmas.
-Lemma 2 (Eigenvalue inequality [32, Corollary 4.3.12]).
+Lemma 2 (Eigenvalue inequality [33, Corollary 4.3.12]).
 For a positive definite symmetric matrix $\mathbf{B}\in M_{N}$ and $c\geq 0$, we have
@@
 with the equality for some $i$ if and only if $c=0$.
-Lemma 3 (Trace inequality [32, Equation 2.4.2.1]).
+Lemma 3 (Trace inequality [33, Equation 2.4.2.1]).
 For a positive definite symmetric matrix $\mathbf{B}\in M_{N}$ and $c\geq 0$, we have